### v2

# Many (most?) column subset selection criteria

Abstract
We consider a variety of criteria for selecting $k$ representative columns from a real $m\times n$ matrix $\boldsymbol{A}$, when sufficiently few columns are required, i.e., $1\leq k\leq\min\{\operatorname{\mathrm{rank}}(\boldsymbol{A}),m/3\}$. The criteria include the following optimization problems: absolute volume and S-optimality maximization; norm, pseudo-inverse norm, and condition minimization number in the two-norm, Frobenius norm and Schatten $p$-norms for $p>2$; stable rank maximization; and the new criterion of relative volume maximization, which is inversely proportional to a power of the condition number. We show that these criteria are NP hard and many do not admit polynomial time approximation schemes (PTAS). To formulate the optimization problems as decision problems, we derive optimal values for the subset selection criteria, as well as expressions for partitioned pseudo-inverses. The results for minimization of the pseudo-inverse in the Frobenius norm are applicable to trace optimization in A-optimal design.
keywords:
volume, condition number, Schatten p-norm, stable rank, NP hard, NP complete, PTAS
[…]
organization = Department of Mathematics, North Carolina State University, city = Raleigh, NC, country = USA
[…]
organization = Department of Mathematics, North Carolina State University, city = Raleigh, NC, country = USA
[…]
abstract
1 Introduction
Given a general [^2] We make no assumptions on the dimensions $m$ and $n$ of $\boldsymbol{A}$. matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$, what does it mean to find $k$ representative columns $\boldsymbol{C}\in\mathbb{R}^{m\times k}$? This problem is known as ‘column subset selection’. Our goal is to establish the complexity classes for the popular subset selection problems in Table 1, in the case when sufficiently few columns are required, i.e., $1\leq k\leq\min\{\operatorname{\mathrm{rank}}(\boldsymbol{A}),m/3\}$.
We show that most criteria in Table 1 are NP hard, which means they cannot be solved in polynomial time unless $P=NP$. Although most subset selection criteria are NP hard, not all are. For example, a submatrix $\boldsymbol{C}$ with minimal Frobenius norm can be computed in polynomial time [^3] A greedy algorithm picks the columns of $\boldsymbol{A}$ with smallest two-norm. (Section 4).
Our strategy is the following: Instead of considering the optimization problem, e.g., find a submatrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with maximal volume, we show NP hardness instead for the associated decision problem, e.g., given $b>0$, find a submatrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ whose volume exceeds $b$. To connect the optimization and decision problems, we derive optimal values for the subset selection criteria (Section 8), so that the optimization problems can be formulated as decision problems for the optimal values.
Having established that a criterion is NP hard, it is reasonable to ask if it can be approximated in polynomial time. This leads to the concept of a Polynomial Time Approximation Scheme (PTAS). A PTAS is a polynomial time algorithm such that for each $\epsilon>0$ there exists a $(1+\epsilon)$-approximation algorithm for minimization problems, and a $(1-\epsilon)$-approximation algorithm for maximization problems [38, Section 1], [43, Definition 1.2]. We show that most criteria in Table 1 do not admit a PTAS either.
1.1 Complexity Theory
We focus on decision problems with a binary output ‘yes’ or ‘no’, and adopt the Turing model of computation, where inputs are rational numbers, and the input size is measured in terms of bits. The only irrational number used in our reduction is $1/\sqrt{3}$, which can also be expressed in terms of a finite number of bit [24, Section 9].
The complexity class $P$ consists of decision problems that can be solved by a deterministic Turing machine in polynomial time, while the complexity class $NP$ consists of decision problems for which a deterministic Turing machine can verify an output of ‘yes’ or ‘no’ in polynomial time. A problem is NP hard if there is polynomial time reduction to this problem from any problem in NP. A decision problem is NP complete if it is in NP and it is also NP hard. Thus, an NP hard problem is at least as hard as an NP complete problem [20, Section 5.1].
A standard strategy for establishing NP hardness performs a polynomial time reduction from an NP complete problem to the problem at hand. This reduction makes it possible to convert any algorithm for solving the problem at hand into a corresponding algorithm for an NP complete problem. In particular, a reduction that operates in polynomial time makes it possible to convert any polynomial time for the problem at hand to a corresponding polynomial time algorithm for the NP complete problem [20, Section 1.5]. Following [9], we choose the NP complete problem ‘Exact cover by 3 sets (X3C)’ [20, Section A.3.1], which appears in Karp’s 1972 list of NP complete problems [28, Section 4].
To show that a subset selection criterion does not admit a PTAS, we create a “gap” between the optimal value and a particular suboptimal value, the so-called inapproximability threshold, and show that it is NP hard to approximate the criterion beyond the threshold, that is, for any value inside the gap.
Unfortunately, creating such a gap is not always feasible [38, Section 1], and we were not able to do so for a few criteria, including Schatten $p$-norm minimization for $p>2$ (Section 4) and condition number minimization in Schatten $p$-norms for $p>2$ (Section 6).
Quantity Criterion NP hard No PTAS Definition Volume $\max_{\boldsymbol{C}}{\operatorname{\mathrm{vol}}(\boldsymbol{C})}$ [9, Theorem 4] [9, Theorem 6](1) $\operatorname{\mathrm{vol}}(\boldsymbol{C})\equiv\prod_{j=1}^{k}{\sigma_{j}(\boldsymbol{C})}$ [18, Theorem 8] Theorem 3 Theorem 4 $\max_{\boldsymbol{C}}{\operatorname{\mathrm{rvol}}(\boldsymbol{C})}$ Theorem 1 Theorem 2 $\operatorname{\mathrm{rvol}}(\boldsymbol{C})\equiv\operatorname{\mathrm{vol}}(\boldsymbol{C})/\|\boldsymbol{C}\|_{2}^{k}$ $\max_{\boldsymbol{C}}{\operatorname{\mathrm{s_{opt}}}(\boldsymbol{C})}$ Theorem 3 Theorem 5 $\operatorname{\mathrm{s_{opt}}}(\boldsymbol{C})\equiv(\operatorname{\mathrm{vol}}(\boldsymbol{C})/\prod_{i=1}^{k}{\|\boldsymbol{C}\boldsymbol{e}_{i}\|_{2})^{1/k}}$ Norm $\min_{\boldsymbol{C}}{\|\boldsymbol{C}\|_{2}}$ [9, Theorem 4] [9, Theorem 5] Theorem 6 Theorem 7 $\min_{\boldsymbol{C}}{\|\boldsymbol{C}\|_{F}}$ No N/A $\min_{\boldsymbol{C}}{\|\boldsymbol{C}\|_{(p)}}$ Theorem 6 Pseudo-inverse $\min_{\boldsymbol{C}}{\|\boldsymbol{C}^{\dagger}\|_{2}}$ [9, Theorem 4] [9, Theorem 7] Theorem 8 Theorem 9 $\min_{\boldsymbol{C}}{\|\boldsymbol{C}^{\dagger}\|_{F}}$ Theorem 8 Theorem 10 $\min_{\boldsymbol{C}}{\|\boldsymbol{C}^{\dagger}\|_{(p)}}$ Theorem 8 Theorem 11 Condition no $\min_{\boldsymbol{C}}{\kappa_{2}(\boldsymbol{C})}$ [9, Theorem 4] [9, Theorem 8] $\kappa_{2}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{2}\|\boldsymbol{C}^{\dagger}\|_{2}$ Theorem 12 Theorem 13 $\min_{\boldsymbol{C}}{\kappa_{F}(\boldsymbol{C})}$ Theorem 12 Theorem 14 $\kappa_{F}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{F}\|\boldsymbol{C}^{\dagger}\|_{F}$ $\min_{\boldsymbol{C}}{\kappa_{D}(\boldsymbol{C})}$ Theorem 12 Theorem 15 $\kappa_{D}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{F}\|\boldsymbol{C}^{\dagger}\|_{2}$ $\min_{\boldsymbol{C}}{\kappa_{D,p}(\boldsymbol{C})}$ Theorem 12 $\kappa_{D,p}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{(p)}\|\boldsymbol{C}^{\dagger}\|_{2}$ Stable rank $\max_{\boldsymbol{C}}{\operatorname{\mathrm{sr}}(\boldsymbol{C})}$ Theorem 16 Theorem 17 $\operatorname{\mathrm{sr}}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{F}^{2}/\|\boldsymbol{C}\|_{2}^{2}$ $\max_{\boldsymbol{C}}{\operatorname{\mathrm{sr}}_{(p)}(\boldsymbol{C})}$ Theorem 16 $\operatorname{\mathrm{sr}}_{(p)}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{(p)}^{p}/\|\boldsymbol{C}\|_{2}^{p}$ Residual $\min_{\boldsymbol{C}}{\rho_{2}(\boldsymbol{C})}$ $\rho_{2}(\boldsymbol{C})\equiv\|(\boldsymbol{I}-\boldsymbol{C}\boldsymbol{C}^{\dagger})\boldsymbol{A}\|_{2}$ $\min_{\boldsymbol{C}}{\rho_{F}(\boldsymbol{C})}$ [40, Theorem 2.2] [6, Theorem 1.1](2) $\rho_{F}(\boldsymbol{C})\equiv\|(\boldsymbol{I}-\boldsymbol{C}\boldsymbol{C}^{\dagger})\boldsymbol{A}\|_{F}$
Table 1: Complexity of criteria for selecting $k\leq\min\{\operatorname{\mathrm{rank}}(\boldsymbol{A}),m/3\}$ representative columns $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ from $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$: (1) Exponential inapproximability of the volume is shown in [10, Theorem 5]. (2) Minimizing the Frobenius norm residual is NP complete [40, Theorem 2.2], and is PTAS inapproximable conditional on the Unique Games Conjecture being true [6, Theorem 1.1].
1.2 Overview
We derive NP hardness results and PTAS approximation thresholds for: relative volume maximization (Section 2); volume and S-optimality maximization (Section 3); two-norm, Frobenius norm and Schatten $p$-norm minimization (Section 4); two-norm, Frobenius norm and Schatten $p$-norm minimization of pseudo-inverses (Section 5); condition number minimization (Section 6); and stable rank maximization (Section 7). To formulate the subset selection criteria as decision problems, we derive optimal values for the criteria (Section 8), as well as expressions for partitioned pseudo-inverses (Section 9).
1.3 Contributions
Our main contributions include the following:
  1. 1.
    We introduce the criterion of ‘relative volume maximization’ and show that it is NP-hard and does not admit a PTAS. Unlike the volume, the relative volume is able to detect whether a matrix is ill-conditioned, because it is inversely proportional to a power of the condition number (Section 2).
  2. 2.
    We present rigorous proofs to derive PTAS inapproximability thresholds that differ from those in [9] for two-norm minimization of the pseudo-inverse (Section 5.2) and two-norm condition number minimization (Section 6.2).
  3. 3.
    We show that the following criteria are NP-hard: $p$-norm pseudo-inverse minimization, $p$-norm condition number minimization, and mixed $p$-norm condition number minimization for $p\geq 2$ (Section 6.1), as well as stable rank maximization and $p$-stable rank maximization for $p>2$ (Section 7.1).
  4. 4.
    We show that the following criteria do not admit a PTAS: Frobenius norm minimization of the pseudo-inverse (Section 5.3) and $p$-norm minimization of the pseudo-inverse for $p>2$ (Section 5.4); Frobenius norm condition number minimization (Section 6.3) and mixed condition number minimization (Section 6.4); and stable rank maximization (Section 7.2).
  5. 5.
    We derive optimal values for the subset selection criteria (Section 8), so that their optimization versions can be formulated as decision problems.
  6. 6.
    We derive expressions and $p$-norm bounds for partitioned pseudo-inverses (Section 9).
1.4 Roadmap
We give a brief sketch of our approaches for deriving NP hardness results (Section 1.4.1) and PTAS results (Section 1.4.2) for the above subset selection criteria.
1.4.1 Extension of the NP-hardness results from [9]
We perform the following steps.
  1. 1.
    Polynomial time reduction of the subset selection criteria to X3C by representing the collection of subsets as a matrix $\boldsymbol{A}$ with unit norm columns.
  2. 2.
    For such matrices with unit-norm columns, we establish optimal values for the subset selection criteria (Section 8), and show that the optimal values are achieved by matrices with orthonormal columns.
    Optimal values are derived for the criteria: maximal volume and S-optimality (Lemma 1), maximal relative volume (Lemma 2), minimal norms (Lemma 4), minimal pseudo-inverse norms (Lemma 5), minimal condition numbers (Lemma 6), and maximal stable ranks (Lemma 8).
  3. 3.
    According to [9], X3C is true is and only if $\boldsymbol{A}$ has a submatrix with orthonormal columns.
  4. 4.
    Combining the two previous items implies that the criterion is optimal if and only if X3C is true.
    This shows NP hardness of the criteria: relative volume maximization (Theorem 1), volume and S-optimality maximization (Theorem 3), norm minimization (Theorem 6), pseudo-inverse norm minimization (Theorem 8), condition number minimization (Theorem 12), and stable rank maximization (Theorem 16).
1.4.2 Systematic approach for extending PTAS results from [9]
We perform the following steps.
  1. 1.
    Monotonicity: Singular value interlacing implies that deleting columns from a matrix can only increase its relative volume (Lemma 3) and decrease the condition number (Lemma 7). In other words, adding columns can only decrease the relative volume and increase the condition number.
  2. 2.
    It therefore suffices to show the lack of a PTAS by establishing approximability thresholds for matrices with 2 columns.
  3. 3.
    Monotonicity implies that matrices with more columns cannot cross the approximability thresholds.
    This shows that the following criteria do not admit a PTAS: maximization of relative volume (Theorem 2), volume (Theorem 4) and S-optimality (Theorem 5); two-norm minimization (Theorem 7); pseudo-inverse minimization in the two- (Theorem 9), Frobenius (Theorem 10) and Schatten $p$-norms (Theorem 11); condition number minimization in the two- (Theorem 13), Frobenius (Theorem 14) and mixed norms (Theorem 15); and maximization of the stable rank (Theorem 17).
1.5 Literature
We review existing work on column subset selection, focused on computational complexity. Early surveys of column subset selection algorithms include [8, 25].
Norm minimization
NP hardness and PTAS inapproximability for the two-norm are shown in [9, Theorems 5 and 7].
Pseudo-inverse norm minimization
NP hardness and PTAS inapproximability are shown in [9, Theorems 4 and 7] for the two-norm. Randomized and deterministic algorithms for wide matrices $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with $m<n$ are given in [1, 2, 29, 34, 35, 37].
In the context of A-optimal design, trace minimization corresponds to minimization of the pseudo-inverse in the Frobenius norm [34, 35]. For instance, if $A$ has full row rank then $\boldsymbol{A}^{\dagger}=\boldsymbol{A}^{T}(\boldsymbol{A}\boldsymbol{A}^{T})^{-1}$ and
$\|\boldsymbol{A}^{\dagger}\|_{F}^{2}=\operatorname{\mathrm{trace}}\left((\boldsymbol{A}^{\dagger})^{T}\boldsymbol{A}^{\dagger}\right)=\operatorname{\mathrm{trace}}\left((\boldsymbol{A}\boldsymbol{A}^{T})^{-1}\right).$
The reduction for $m=k$ is based on Odd Cycle Packing, which consists of finding a maximal family of vertex-disjoint odd cycles in a simple undirected graph.
Condition number minimization
NP hardness and PTAS inapproximability are shown in [9, Theorems 4 and 8] for two-norm condition numbers.
Volume maximization
NP hardness and PTAS inapproximability are shown in [9, Theorems 4 and 6], while the exponential inapproximability of the volume is presented in [10, Theorem 5], [18, Theorem 8]. The latter involves a reduction based on Odd Cycle Packing for matrices $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{A})=m$, $k=m$, $m=1320d$ and $n=2112d$ for some $d>0$.
An early result [42, Theorem 3] in the context of D-optimality shows the NP-hardness of volume maximization via a reduction from Hamiltonian Circuit, where $k\geq n$ rows $\boldsymbol{C}\in\mathbb{R}^{k\times n}$ from an $m\times n$ adjacency matrix are to be selected to maximize $\operatorname{\mathrm{vol}}(\boldsymbol{C})^{2}=\det(\boldsymbol{C}^{T}\boldsymbol{C})$.
Greedy algorithms are presented in [9, 14], and local deterministic and randomized algorithms in [22, 23, 33, 36]. This is extended to Hilbert spaces in [30]. The importance of local volume maximization for revealing the rank is demonstrated in [13].
The S-optimality criterion bears a resemblance to the relative volume and was introduced in [39, (3.5)] and [31, (3.10)].
Residual Minimization
The NP completeness in the Frobenius norm is established in [40, Theorem 2.2], and PTAS inapproximability conditional on the Unique Games Conjecture being true [6, Theorem 1.1]. Algorithms for the two- and Frobenius norms are given in [5].
Early randomized algorithms, based on volume sampling are given in [5, 16, 17]. A statistical perspective is provided in [41], and it is shown that subset selection is equivalent to the method of principal variables.
1.6 Assumptions
Let $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ be a general matrix with $\operatorname{\mathrm{rank}}(\boldsymbol{A})\geq k$, and singular values $\sigma_{1}(\boldsymbol{A})\geq\cdots\geq\sigma_{\min\{m,n\}}(\boldsymbol{A})\geq 0$. Let $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ be a submatrix of $k$ columns of $\boldsymbol{A}$, with singular values
$\sigma_{1}(\boldsymbol{C})\geq\cdots\geq\sigma_{k}(\boldsymbol{C})\geq 0.$ (1)
Singular value interlacing [26, 7.3.P44] implies
$\sigma_{\min\{m,n\}-k+j}(\boldsymbol{A})\leq\sigma_{j}(\boldsymbol{C})\leq\sigma_{j}(\boldsymbol{A}),\qquad 1\leq j\leq k.$ (2)
We denote by $\boldsymbol{I}_{n}\equiv\begin{bmatrix}\boldsymbol{e}_{1}&\cdots&\boldsymbol{e}_{n}\end{bmatrix}\in{\mathbb{R}}^{n\times n}$ the identity matrix, and by $\boldsymbol{e}_{j}\in\mathbb{R}^{n}$ its columns.
2 Relative Volume
We introduce the concept of maximal volume, and show that relative volume maximization is NP hard (Section 2.1) and admits no PTAS (Section 2.2).
The relative volume of a matrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$, defined as
$\operatorname{\mathrm{rvol}}(\boldsymbol{C})\equiv\frac{\operatorname{\mathrm{vol}}(\boldsymbol{C})}{\|\boldsymbol{C}\|_{2}^{k}}=\prod_{j=1}^{k}{\frac{\sigma_{j}(\boldsymbol{C})}{\sigma_{1}(\boldsymbol{C})}},$
where the volume of a matrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$ is defined as
$\operatorname{\mathrm{vol}}(\boldsymbol{C})\equiv\sqrt{\det(\boldsymbol{C}^{T}\boldsymbol{C})}=\prod_{j=1}^{k}{\sigma_{j}(\boldsymbol{C})}.$ (3)
If $\boldsymbol{C}$ is square, then $\operatorname{\mathrm{vol}}(\boldsymbol{C})=|\det(\boldsymbol{C})|$.
Geometrically, the relative volume is the ratio of the parallelepiped volume to the volume of the cube that contains the parallelepiped.
Example
In contrast to the volume, the relative volume has the advantage of being able to detect when a matrix is ill-conditioned. For instance, the matrix
$\boldsymbol{A}=\begin{bmatrix}1/\epsilon&0\\ 0&\epsilon\end{bmatrix}\qquad\text{where}\quad 0<\epsilon\ll 1$
is ill-conditioned with condition number $\kappa_{2}(\boldsymbol{A})=1/\epsilon^{2}$. However, the volume does not detect this, since $\operatorname{\mathrm{vol}}(\boldsymbol{A})=1$, while the relative volume does, $\operatorname{\mathrm{rvol}}(\boldsymbol{A})=\epsilon^{2}=1/\kappa_{2}(\boldsymbol{C})$.
Remark 2.1
The relative volume is inversely proportional to some power of the two-norm condition number. Let $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$. Then
$\left(\frac{1}{\kappa_{2}(\boldsymbol{C})}\right)^{k-1}\leq\operatorname{\mathrm{rvol}}(\boldsymbol{C})\leq\frac{1}{\kappa_{2}(\boldsymbol{C})}.$
Next, we want to find $k$ columns $\boldsymbol{C}$ of $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with maximal relative volume, $\max_{\boldsymbol{C}}{\operatorname{\mathrm{rvol}}(\boldsymbol{C})}$.
2.1 Decision problem
Following [20], we state the decision problem for the maximal relative volume as follows:
Instance: Matrix $\boldsymbol{A}\in\mathbb{R}^{m\times n}$ with unit-norm columns $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$; integer $1\leq k\leq\min\{\operatorname{\mathrm{rank}}(\boldsymbol{A}),m/3\}$; and parameter $0<b\leq 1$.
Question: Is there a column submatrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ of $\boldsymbol{A}$ such that
$\operatorname{\mathrm{rvol}}(\boldsymbol{C})\geq b.$ (4)
Theorem 1
The decision problem of relative volume maximization is NP hard.
Proof 1
Following the proof of [9, Proof of Theorem 4], we perform a polynomial-time reduction from the NP-complete problem Exact Cover by 3-Sets (X3C) [20, Section A.3.1] to relative volume maximization.
Instance: Set $\mathcal{S}=\{1,\dots,3M\}$ and collection of sets $\mathcal{K}=\{\mathcal{K}_{1},\dots,\mathcal{K}_{n}\}$, where each $\mathcal{K}_{i}\subset\mathcal{S}$ has cardinality $3$.
Question: Is there a subset $\mathcal{K}^{\prime}\subset\mathcal{K}$ that forms an exact cover [^4] An exact cover means that every element of $\mathcal{S}$ appears exactly once in $\mathcal{K}^{\prime}$. for $\mathcal{S}$?
Assume that $\mathcal{S}=\{1,\dots,3M\}$ and a collection of distinct subsets $\mathcal{K}=\{\mathcal{K}_{1},\dots,\mathcal{K}_{n}\}$ represent an instance of X3C. We show that this instance can be solved if we can construct a matrix $\boldsymbol{A}$ with a submatrix $\boldsymbol{C}_{*}$ that solves (4).
To this end, let $\boldsymbol{A}\in\mathbb{R}^{(3M)\times n}$ be a matrix where each column corresponds to a 3-element set $\mathcal{K}_{j}$, and the non-zero elements in the column represent the elements of $\mathcal{K}_{j}$,
$a_{ij}=\begin{cases}\frac{1}{\sqrt{3}}&\text{if $i\in\mathcal{K}_{j}$}\\ 0&\text{otherwise}\end{cases},\qquad 1\leq i\leq 3M,\ 1\leq j\leq n.$ (5)
Each column of $\boldsymbol{A}$ has exactly 3 non-zero elements and $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$. In the reduced instance we set $m=3M$, $k=M$, and $b=1$. To construct matrices $\boldsymbol{A}$ with more than $3k$ rows, just append the appropriate number of zero rows. This does not change the singular values of $\boldsymbol{A}$. Clearly, constructing $\boldsymbol{A}$ requires polynomial time.
According to [9, Proof of Theorem 4], the instance of X3C is true if and only if there exists a submatrix $\boldsymbol{C}_{*}\in\mathbb{R}^{(3M)\times M}$ with orthogonal columns. Lemma 2 in Section 8.2 implies that this is equivalent to $\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{*})=1$. Due to this equivalence and the structure of $\boldsymbol{A}$, determining whether $\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{*})=1$ simply amounts to finding orthogonal columns based on the sparsity pattern of $\boldsymbol{C}_{*}$. Therefore, the instance of X3C is true if and only if $\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{*})=1$. Thus, we have constructed a polynomial time reduction from an NP complete problem, establishing that relative volume maximization is NP hard.
2.2 No PTAS for relative volume maximization
Lemma 2 implies that the maximal relative volume equals 1. We derive an inapproximability threshold of $\sqrt{1/2}$.
Theorem 2
It is NP hard to approximate relative volume maximization within a factor of $1/\sqrt{2}$. Thus, unless P=NP, there is no PTAS for relative volume maximization.
Proof 2
This is analogous to the proof of [9, Theorem 5] for volume maximization. We derive an upper bound for $\operatorname{\mathrm{rvol}}(\boldsymbol{C})$ for the reduced instance of X3C when it is false. Lemma 3 in Section 8.2 implies that the relative volume increases with the removal of columns. Hence, the largest relative volume is determined by finding two columns for a false instance of X3C.
Let $\boldsymbol{A}$ be constructed as in (5), and assume that the X3C instance is false. Then any collection $\mathcal{K}$ of cardinality $k=M$ has at least two sets $\mathcal{K}_{i}$ and $\mathcal{K}_{j}$ with $i\neq j$ and a non-empty intersection. Since the subsets $\mathcal{K}_{j}$ are distinct, the overlap consists of 1 or of 2 elements. Let $\boldsymbol{C}\in\mathbb{R}^{(3M)\times M}$ be a column submatrix of $\boldsymbol{A}\in\mathbb{R}^{(3M)\times n}$ that represents such a collection.
  1. 1.
    Consider the case where the two sets share a single element, that is, $|\mathcal{K}_{i}\cap\mathcal{K}_{j}|=1$. Let the sets be represented by columns $\boldsymbol{c}_{i}\equiv\boldsymbol{C}\boldsymbol{e}_{i}$ and $\boldsymbol{c}_{j}\equiv\boldsymbol{C}\boldsymbol{e}_{j}$. The non-zero elements of the submatrix $\hat{\boldsymbol{C}}\equiv\begin{bmatrix}\boldsymbol{c}_{i}&\boldsymbol{c}_{j}\end{bmatrix}$ are, up to row permutations,
    $\begin{bmatrix}1/\sqrt{3}&1/\sqrt{3}&1/\sqrt{3}&0&0\\ 0&0&1/\sqrt{3}&1/\sqrt{3}&1/\sqrt{3}\end{bmatrix}^{T}\in\mathbb{R}^{5\times 2}.$ (6)
    With appropriate orthogonal row rotations this matrix can be further reduced to
    $\begin{bmatrix}0&1&0&0&0\\ 0&1/3&\sqrt{8}/3&0&0\end{bmatrix}^{T}.$
    Thus $\hat{\boldsymbol{C}}$ has the same singular values as the matrix
    $\begin{bmatrix}1&1/3\\ 0&\sqrt{8}/3\end{bmatrix}$,
    whose singular values are $2/\sqrt{3}$ and $\sqrt{2/3}$. Hence, $\sigma_{1}(\hat{\boldsymbol{C}})=2/\sqrt{3}$ and $\sigma_{2}(\hat{\boldsymbol{C}})=\sqrt{2/3}$. Lemma 3 in Section 8.2 implies
    $\operatorname{\mathrm{rvol}}(\boldsymbol{C})\leq\operatorname{\mathrm{rvol}}(\hat{\boldsymbol{C}})=\sigma_{2}(\hat{\boldsymbol{C}})/\sigma_{1}(\hat{\boldsymbol{C}})=1/\sqrt{2}.$
  2. 2.
    Now consider the case where the sets share two elements. That is, if $|\mathcal{K}_{i}\cap\mathcal{K}_{j}|=2$, we can find two columns $\hat{\boldsymbol{C}}\equiv\begin{bmatrix}\boldsymbol{c}_{i}&\boldsymbol{c}_{j}\end{bmatrix}$ of $\boldsymbol{C}$ whose non-zero elements are, up to row permutations,
    $\begin{bmatrix}1/\sqrt{3}&1/\sqrt{3}&1/\sqrt{3}&0\\ 0&1/\sqrt{3}&1/\sqrt{3}&1/\sqrt{3}\end{bmatrix}^{T}\in\mathbb{R}^{4\times 2}.$ (7)
    After row rotations, the nonzero elements are
    $\begin{bmatrix}1&2/3\\ 0&\sqrt{5}/3\end{bmatrix}$,
    whose singular values are $\sqrt{5/3}$ and $\sqrt{1/3}$. Hence, $\sigma_{1}(\hat{\boldsymbol{C}})=\sqrt{5/3}$ and $\sigma_{2}(\hat{\boldsymbol{C}})=\sqrt{1/3}$. Lemma 3 in Section 8.2 implies
    $\operatorname{\mathrm{rvol}}(\boldsymbol{C})\leq\operatorname{\mathrm{rvol}}(\hat{\boldsymbol{C}})=\sigma_{2}(\hat{\boldsymbol{C}})/\sigma_{1}(\hat{\boldsymbol{C}})=1/\sqrt{5}.$
    Thus, the relative volume of (7) is smaller than the one of (6).
Considering the overlap of only two sets is sufficient for the following reason. Suppose that $\ell<k$ columns of $\boldsymbol{C}$ overlap, and they are represented by the submatrix $\boldsymbol{C}_{\ell}$. Lemma 3 in Section 8.2 implies that $\operatorname{\mathrm{rvol}}(\boldsymbol{C})\leq\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{\ell})$. Let $\boldsymbol{C}_{2}$ be any 2-column submatrix of $\boldsymbol{C}_{\ell}$. Again, Lemma 3 implies $\operatorname{\mathrm{rvol}}(\boldsymbol{C})\leq\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{\ell})\leq\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{2})$. The non-zero elements of $\boldsymbol{C}_{2}$ can only be of the form (6) or (7). Between the two, (6) achieves the largest relative volume $1/\sqrt{2}$, which implies $\operatorname{\mathrm{rvol}}(\boldsymbol{C})\leq 1/\sqrt{2}$.
Summary: As a consequence, it is NP hard to approximate the relative volume by a factor of $1/\sqrt{2}$. If an algorithm for relative volume maximization did indeed return a set of columns $\boldsymbol{C}$ with $\operatorname{\mathrm{rvol}}(\boldsymbol{C})>1/\sqrt{2}$, then this instance of X3C would be true. Thus, relative volume maximization could distinguish between true and false instances of X3C, implying $P=NP$.
3 Volume and S-optimality
We show that volume and S-optimality maximization are NP hard (Section 3.1), give a rigorous proof of the PTAS inapproximability threshold for volume maximization (Section 3.2), and show that S-optimality maximization does not admit a PTAS (Section 3.3).
We want to find $k$ columns $\boldsymbol{C}$ of $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with maximal volume, $\max_{\boldsymbol{C}}{\operatorname{\mathrm{vol}}(\boldsymbol{C})}$. This criterion is NP hard and does not admit a PTAS [9]. The best deterministic algorithm [23] is a local volume maximization that produces a matrix $\boldsymbol{C}$ whose singular values are maximal subject to pairwise permutations of columns.
The S-optimality of a matrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$ is defined as [39, (3.5)] and [31, (3.10)]
$\operatorname{\mathrm{s_{opt}}}(\boldsymbol{C})\equiv\left(\frac{\operatorname{\mathrm{vol}}(\boldsymbol{C})}{\prod_{i=1}^{k}{\|\boldsymbol{C}\boldsymbol{e}_{i}\|_{2}}}\right)^{1/k}=\left(\prod_{i=1}^{k}{\frac{\sigma_{i}(\boldsymbol{C})}{\|\boldsymbol{C}\boldsymbol{e}_{i}\|_{2}}}\right)^{1/k},$
and greedy algorithms are presented for maximizing the S-optimality. We want to find $k$ columns $\boldsymbol{C}$ of $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with maximal S-optimality, $\max_{\boldsymbol{C}}{\operatorname{\mathrm{s_{opt}}}(\boldsymbol{C})}$.
3.1 Decision problems
Given a matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with unit-norm columns $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$; integer $1\leq k\leq\min\{\operatorname{\mathrm{rank}}(\boldsymbol{A}),m/3\}$; and parameter $0\leq b\leq 1$.
  1. 1.
    Volume maximization:
    Does $\boldsymbol{A}$ have a submatrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{vol}}(\boldsymbol{C})\geq b$?
  2. 2.
    S-optimality maximization:
    Does $\boldsymbol{A}$ have a submatrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{s_{opt}}}(\boldsymbol{C})\geq b$?
The NP hardness of volume maximization was already established in [9, Theorem 4].
Theorem 3
The decision problems: volume maximization and S-optimality maximization are NP hard.
Proof 3
The idea is that the reduction to X3C in Theorem 1 only requires orthonormality: The instance of X3C is true if and only if $\boldsymbol{A}$ has a submatrix $\boldsymbol{C}\in\mathbb{R}^{(3M)\times M}$ with orthonormal columns.
As in (5), we construct the matrix $\boldsymbol{A}\in\mathbb{R}^{3M\times n}$ where each column has exactly 3 non-zero elements and $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$. According to [9, Proof of Theorem 4], the instance of X3C is true if and only if $\boldsymbol{A}$ has a submatrix $\boldsymbol{C}\in\mathbb{R}^{(3M)\times M}$ with orthonormal columns. Lemma 1 in Section 8.1 implies that this is equivalent to $\boldsymbol{C}$ satisfying the equalities $\operatorname{\mathrm{vol}}(\boldsymbol{C})=1$ or $\operatorname{\mathrm{s_{opt}}}(\boldsymbol{C})=1$. Thus, volume maximization and S-optimality maximization are NP hard.
3.2 No PTAS for volume maximization
Lemma 1 in Section 8.1 implies that a matrix with unit-norm columns has maximal volume equal to 1. We derive an inapproximability threshold of $2\sqrt{2}/3$.
Theorem 4
It is NP hard to approximate volume maximization within a factor of $2\sqrt{2}/3$. Thus, unless P=NP, there is no PTAS for volume maximization.
Proof 4
The setup is the same as in the proof of Theorem 2. Assume that the X3C instance is false.
The proof is by induction over the number of columns $k$.
  • 1.
    For $k=2$, two distinct sets can overlap in one or in two elements. When the sets share a single element, the nonzero elements (6) of the relevant two columns $\hat{\boldsymbol{C}}_{2}$ have singular values $\sigma_{1}(\hat{\boldsymbol{C}}_{2})=2/\sqrt{3}$ and $\sigma_{2}(\hat{\boldsymbol{C}}_{2})=\sqrt{3/2}$. When the sets share two elements, the nonzero elements (7) of the relevant two columns $\hat{\boldsymbol{C}}_{2}$ have singular values $\sigma_{1}(\hat{\boldsymbol{C}}_{2})=\sqrt{5/3}$ and $\sigma_{2}(\hat{\boldsymbol{C}}_{2})=\sqrt{1/3}$. Between the two, (6) achieves the largest volume. Hence, for matrices $\hat{\boldsymbol{C}}_{2}$ with $k=2$ columns, (6) implies that
    $\operatorname{\mathrm{vol}}(\hat{\boldsymbol{C}}_{2})=\sigma_{1}(\hat{\boldsymbol{C}}_{2})\sigma_{2}(\hat{\boldsymbol{C}}_{2})=\tfrac{2}{\sqrt{3}}\sqrt{\tfrac{2}{3}}=\tfrac{2\sqrt{2}}{3}.$
  • 2.
    For matrices $\hat{\boldsymbol{C}}_{k-1}$ with k1k-1 columns, assume that
    $\operatorname{\mathrm{vol}}(\hat{\boldsymbol{C}}_{k-1})\leq\tfrac{2\sqrt{2}}{3}.$
  • 3.
    Let $\hat{\boldsymbol{C}}_{k}\equiv\begin{bmatrix}\hat{\boldsymbol{C}}_{k-1}&\boldsymbol{c}\end{bmatrix}$ be a matrix with $k$ columns, where $\|\boldsymbol{c}\|_{2}=1$. The Gram matrix is
    $\hat{\boldsymbol{C}}_{k}^{T}\hat{\boldsymbol{C}}_{k}=\begin{bmatrix}\hat{\boldsymbol{C}}_{k-1}^{T}\hat{\boldsymbol{C}}_{k-1}&\hat{\boldsymbol{C}}_{k-1}^{T}\boldsymbol{c}\\ \boldsymbol{c}^{T}\hat{\boldsymbol{C}}_{k-1}&\boldsymbol{c}^{T}\boldsymbol{c}\end{bmatrix}$
    and [12, (2)] implies
    $\operatorname{\mathrm{vol}}(\hat{\boldsymbol{C}}_{k})^{2}=\det(\hat{\boldsymbol{C}}_{k-1}^{T}\hat{\boldsymbol{C}}_{k-1})\det(\boldsymbol{c}^{T}\boldsymbol{c}-\boldsymbol{c}^{T}\hat{\boldsymbol{C}}_{k}^{T}\hat{\boldsymbol{C}}_{k}^{\dagger}\boldsymbol{c}).$ (8)
    Since the second factor is a scalar and contains the orthogonal projector $\boldsymbol{P}\equiv\boldsymbol{I}-\hat{\boldsymbol{C}}_{k}^{T}\hat{\boldsymbol{C}}_{k}^{\dagger}$, we can write the second factor as
    $\det(\boldsymbol{c}^{T}\boldsymbol{c}-\boldsymbol{c}^{T}\hat{\boldsymbol{C}}_{k}^{T}\hat{\boldsymbol{C}}_{k}^{\dagger}\boldsymbol{c})=\boldsymbol{c}^{T}\boldsymbol{P}\boldsymbol{c}=\|\boldsymbol{P}\boldsymbol{c}\|_{2}^{2}\leq\|\boldsymbol{c}\|_{2}^{2}=1.$
    Inserting this into (8) gives
    $\operatorname{\mathrm{vol}}(\hat{\boldsymbol{C}}_{k})^{2}\leq\det(\hat{\boldsymbol{C}}_{k-1}^{T}\hat{\boldsymbol{C}}_{k-1})=\operatorname{\mathrm{vol}}(\hat{\boldsymbol{C}}_{k-1})^{2}\leq\left(\tfrac{2\sqrt{2}}{3}\right)^{2}.$
As in the proof of Theorem 2, we conclude that it is NP hard to approximate the maximal volume within a factor of $2\sqrt{2}/3$. Thus, there is no PTAS unless P=NP.
Theorem 4 presents a rigorous proof for the inapproximability threshold from [9, Theorem 6]. The exponential inapproximability of the volume is presented in [10, Theorem 5], in the sense that there exists $0<\delta<1$ and $c>0$ such that volume maximization is not approximable with $2^{-ck}$ for $k=\delta n$ unless $P=NP$. However, the estimation of the volume from inner products of selected vectors in [10, Proof of Lemma 16] is not clear.
3.3 No PTAS for S-optimality maximization
Lemma 1 in Section 8.1 implies that a matrix with unit-norm columns has maximal S-optimality of $1$. We derive an inapproximability threshold of $(2\sqrt{2}/3)^{1/k}$.
Theorem 5
It is NP hard to approximate S-optimality maximization within a factor of $(2\sqrt{2}/3)^{1/k}$. Thus, unless P=NP, there is no PTAS for S-optimality maximization.
Proof 5
This follows from the assumption of $\boldsymbol{C}$ having unit-norm columns, so that $\operatorname{\mathrm{s_{opt}}}(\boldsymbol{S})=\operatorname{\mathrm{vol}}(\boldsymbol{C})^{1/k}$, and from Theorem 4.
4 Two-norm and Schatten p-norms
We show that minimization of the two-norm and Schatten $p$-norms for $p>2$ is NP hard (Section 4.1), and that there is no PTAS for two-norm minimization (Section 4.2).
For $p\geq 1$, the Schatten $p$-norm of a matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ is defined as
$\|\boldsymbol{A}\|_{(p)}\equiv\left(\sum_{j=1}^{\min\{m,n\}}\sigma_{j}(\boldsymbol{A})^{p}\right)^{1/p}.$ (9)
Special cases include the Frobenius norm $\|\boldsymbol{A}\|_{(2)}=\|\boldsymbol{A}\|_{F}$, and the two-norm $\|\boldsymbol{A}\|_{(\infty)}=\|\boldsymbol{A}\|_{2}$. Here we consider only integer values of $p$. We want to find $k$ columns $\boldsymbol{C}$ of $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with minimal Schatten $p$-norm, $\max_{\boldsymbol{C}}{\|\boldsymbol{C}\|_{(p)}}$.
4.1 Decision problems
Given a matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with unit-norm columns $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$; integer $1\leq k\leq\min\{\operatorname{\mathrm{rank}}(\boldsymbol{A}),m/3\}$; and parameter $b>0$.
  1. 1.
    Two-norm minimization:
    Does $\boldsymbol{A}$ have a submatrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\|\boldsymbol{C}\|_{2}\leq b$?
  2. 2.
    Schatten $p$-norm minimization:
    For $p>2$, does $\boldsymbol{A}$ have a submatrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\|\boldsymbol{C}\|_{(p)}\leq b$?
Theorem 6
The decision problems: two-norm minimization and Schatten $p$-norm minimization are NP hard.
Proof 6
The proof is the same as that of Theorem 3, except we use Lemma 4 in Section 8.3 instead.
Remark 8.1 in Section 8.3 explains the requirement $p>2$: The meaningful bound for $p<2$ is an upper bound instead of a lower bound.
In the case $p=2$, minimizing the Frobenius norm can be accomplished by a polynomial time algorithm that relies on the identity
$\|\boldsymbol{C}\|_{F}^{2}=\sum_{j=1}^{k}\|\boldsymbol{C}\boldsymbol{e}_{j}\|_{2}^{2},$
and constructs $\boldsymbol{C}$ from $k$ columns of $\boldsymbol{A}$ with smallest two-norm.
4.2 No PTAS for two-norm minimization
Lemma 4 in Section 8.3 implies that a matrix with unit-norm columns has a minimal two-norm equal to 1. We give a rigorous derivation of the inapproximability threshold $2/\sqrt{3}$ from [9, Theorem 5].
Theorem 7
It is NP hard to approximate two-norm minimization within a factor of $2/\sqrt{3}$. Thus, unless P=NP, there is no PTAS for two-norm minimization.
Proof 7
The setup is the same as in the proof of Theorem 2, and we present an induction over the number of columns $k$ in $\boldsymbol{C}$ to show that $\|\boldsymbol{C}\|_{2}\geq 2/\sqrt{3}$.
  • 1.
    For $k=2$, two distinct sets can overlap in one or in two elements. When the sets share a single element, the nonzero elements (6) of the relevant two columns $\hat{\boldsymbol{C}}_{2}$ have largest singular value $\|\hat{\boldsymbol{C}}_{2}\|_{2}=\sigma_{1}(\hat{\boldsymbol{C}}_{2})=2/\sqrt{3}$. When the sets share two elements, the nonzero elements (7) of the relevant two columns $\hat{\boldsymbol{C}}_{2}$ have largest singular value $\|\hat{\boldsymbol{C}}_{2}\|_{2}=\sigma_{1}(\hat{\boldsymbol{C}}_{2})=\sqrt{5/3}$. Between the two, (6) represents the smaller of the two large singular values. Hence, for matrices $\hat{\boldsymbol{C}}_{2}$ with $k=2$ columns $\|\hat{\boldsymbol{C}}_{2}\|_{2}=\sigma_{1}(\hat{\boldsymbol{C}}_{2})\geq 2/\sqrt{3}$.
  • 2.
    For matrices $\hat{\boldsymbol{C}}_{k-1}$ with k1k-1 columns, assume that
    $\|\hat{\boldsymbol{C}}_{k-1}\|_{2}=\sigma_{1}(\hat{\boldsymbol{C}}_{k-1})\geq 2/\sqrt{3}.$
  • 3.
    Let $\hat{\boldsymbol{C}}_{k}\equiv\begin{bmatrix}\hat{\boldsymbol{C}}_{k-1}&\boldsymbol{c}\end{bmatrix}$ be a matrix with $k$ columns. From singular value interlacing [21, Corollary 8.6.3] follows $\sigma_{1}(\hat{\boldsymbol{C}}_{k})\geq\sigma_{1}(\hat{\boldsymbol{C}}_{k-1})$. Together with the above induction hypothesis this implies
    $\|\hat{\boldsymbol{C}}_{k}\|_{2}=\sigma_{1}(\hat{\boldsymbol{C}}_{k})\geq\sigma_{1}(\hat{\boldsymbol{C}}_{k-1})\geq 2/\sqrt{3}.$
5 Pseudo-inverse norms
We show that minimizing the Schatten $p$-norms of the pseudo-inverse is NP hard (Section 5.1), derive a new approximability threshold for minimizing the two-norm of the pseudo-inverse (Section 5.2), and show that the Frobenius norm (Section 5.3) and Schatten $p$-norm (Section 5.4) minimization of the pseudo-inverse do not admit a PTAS.
We want to find $k$ columns $\boldsymbol{C}$ of $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ whose pseudo-inverse has minimal Schatten $p$-norm, $\min_{\boldsymbol{C}}{\|\boldsymbol{C}^{\dagger}\|_{(p)}}$.
5.1 Decision problems
Given a matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with unit-norm columns $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$; integer $1\leq k\leq\min\{\operatorname{\mathrm{rank}}(\boldsymbol{A}),m/3\}$; and parameter $b>0$.
  1. 1.
    Two-norm pseudo-inverse minimization:
    Does $\boldsymbol{A}$ have a submatrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\|\boldsymbol{C}^{\dagger}\|_{2}\leq b$?
  2. 2.
    Frobenius-norm pseudo-inverse minimization:
    Does $\boldsymbol{A}$ have a submatrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\|\boldsymbol{C}^{\dagger}\|_{F}\leq b$?
  3. 3.
    Schatten $p$-norm pseudo-inverse minimization:
    Does $\boldsymbol{A}$ have a submatrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\|\boldsymbol{C}^{\dagger}\|_{(p)}\leq b$?
Theorem 8
The decision problems: minimizing the pseudo-inverse norm for the two-norm, Frobenius norm, and Schatten $p$-norms for $p>2$ are NP hard.
Proof 8
The proof for all the norms is the same as that of Theorem 3 but uses Lemma 5 instead.
5.2 No PTAS for two-norm minimization of the pseudo-inverse
Lemma 5 in Section 8.5 implies that the pseudo-inverse of a matrix with unit-norm columns has minimal two-norm of 1. We derive an inapproximability threshold of $\sqrt{3/2}$.
Theorem 9
It is NP hard to approximate two-norm minimization of the pseudo-inverse within a factor of $2/\sqrt{3}$. Thus, unless P=NP, there is no PTAS for two-norm minimization of the pseudo-inverse.
Proof 9
The setup is the same as in the proof of Theorem 2, and we present an induction over the number of columns $k$ in $\boldsymbol{C}$ to show that $\|\boldsymbol{C}^{\dagger}\|_{2}\geq\sqrt{3/2}$.
  • 1.
    For $k=2$, two distinct sets can overlap in one or in two elements. When the sets share a single element, the nonzero elements (6) of the relevant two columns $\hat{\boldsymbol{C}}_{2}$ have smallest singular value $\sigma_{2}(\hat{\boldsymbol{C}}_{2})=\sqrt{2/3}$. When the sets share two elements, the nonzero elements (7) of the relevant two columns $\hat{\boldsymbol{C}}_{2}$ have smallest singular value $\sigma_{2}(\hat{\boldsymbol{C}}_{2})=\sqrt{1/3}$. Between the two, (6) represents the largest of the two small singular values. Hence, for matrices $\hat{\boldsymbol{C}}_{2}$ with two columns $\|\hat{\boldsymbol{C}}_{2}^{\dagger}\|_{2}=1/\sigma_{2}(\hat{\boldsymbol{C}}_{2})\geq\sqrt{3/2}$.
  • 2.
    For matrices $\hat{\boldsymbol{C}}_{k-1}$ with k1k-1 columns, assume that
    $\|\hat{\boldsymbol{C}}_{k-1}^{\dagger}\|_{2}=\sigma_{k-1}(\hat{\boldsymbol{C}}_{k-1})\geq\sqrt{3/2}.$
  • 3.
    Let $\hat{\boldsymbol{C}}_{k}\equiv\begin{bmatrix}\hat{\boldsymbol{C}}_{k-1}&\boldsymbol{c}\end{bmatrix}$ be a matrix with $k$ columns. From singular value interlacing [21, Corollary 8.6.3] follows $\sigma_{k-1}(\hat{\boldsymbol{C}}_{k-1})\geq\sigma_{k}(\hat{\boldsymbol{C}}_{k})$. Together with the above induction hypothesis this implies
    $\|\hat{\boldsymbol{C}}_{k}^{\dagger}\|_{2}=1/\sigma_{k}(\hat{\boldsymbol{C}}_{k})\geq 1/\sigma_{k-1}(\hat{\boldsymbol{C}}_{k-1})\geq\sqrt{3/2}.$
The inapproximability threshold of $\sqrt{3/2}$ in Theorem 9 differs from the one in [9, Theorem 7], which is $(\sqrt{3/2})^{\frac{1}{k-1}}$. The discrepancy seems to be due to the estimation of the smallest singular value from the volume.
5.3 No PTAS for Frobenius norm minimization of the pseudo-inverse
Lemma 5 in Section 8.4 implies that the pseudo-inverse of a matrix with $k$ unit-norm columns has a minimal Frobenius norm of $\sqrt{k}$. We derive an inapproximability threshold of $\sqrt{k+\tfrac{1}{4}}$.
Theorem 10
It is NP hard to approximate Frobenius norm minimization of the pseudo-inverse within a factor of $\sqrt{1+\tfrac{1}{4k}}$. Thus, unless P=NP, there is no PTAS for Frobenius norm minimization of the pseudo-inverse.
Proof 10
The setup is the same as in the proof of Theorem 2. Assume that X3C instance is false, and assume that any two columns of $\boldsymbol{C}$ overlap.
We present induction over the number of columns $k$ to show that $\|\boldsymbol{C}^{\dagger}\|_{F}\geq\sqrt{k+\tfrac{1}{4}}$.
  • 1.
    For $k=2$, two distinct sets can overlap in one or in two elements. When the sets share a single element, the nonzero elements (6) of the relevant two columns $\hat{\boldsymbol{C}}_{2}$ have singular values $\sigma_{1}(\hat{\boldsymbol{C}}_{2})=2/\sqrt{3}$ and $\sigma_{2}(\hat{\boldsymbol{C}}_{2})=\sqrt{3/2}$. When the sets share two elements, the nonzero elements (7) of the relevant two columns $\hat{\boldsymbol{C}}_{2}$ have singular values $\sigma_{1}(\hat{\boldsymbol{C}}_{2})=\sqrt{5/3}$ and $\sigma_{2}(\hat{\boldsymbol{C}}_{2})=\sqrt{1/3}$. Between the two, (6) achieves the smallest of two-norm of the pseudoinverse Hence, for matrices $\hat{\boldsymbol{C}}_{2}$ with $k=2$ columns, (6) implies that
    $\|\hat{\boldsymbol{C}}_{2}\|_{F}^{2}=\left(\tfrac{1}{\sigma_{1}(\hat{\boldsymbol{C}}_{2})}\right)^{2}+\left(\tfrac{1}{\sigma_{2}(\hat{\boldsymbol{C}}_{2})}\right)^{2}=\left(\tfrac{\sqrt{3}}{2}\right)^{2}+\left(\sqrt{\tfrac{3}{2}}\right)^{2}=2+\tfrac{1}{4}=k+\tfrac{1}{4}.$
  • 2.
    For matrices $\hat{\boldsymbol{C}}_{k-1}$ with k1k-1 columns, assume that
    $\|\hat{\boldsymbol{C}}_{k-1}^{\dagger}\|_{F}^{2}\geq k-1+\tfrac{1}{4}.$
  • 3.
    Let $\hat{\boldsymbol{C}}_{k}\equiv\begin{bmatrix}\hat{\boldsymbol{C}}_{k-1}&\boldsymbol{c}\end{bmatrix}$ be a matrix with $k$ columns, where $\|\boldsymbol{c}\|_{2}=1$. From Lemma 10 in Section 9 follows
    $\|\hat{\boldsymbol{C}}_{k}^{\dagger}\|_{F}^{2}\geq\|\hat{\boldsymbol{C}}_{k-1}^{\dagger}\|_{F}^{2}+\|\boldsymbol{c}^{\dagger}\|_{2}^{2}\geq\left(k-1+\tfrac{1}{4}\right)+1=k+\tfrac{1}{4}.$
5.4 No PTAS for Schatten p-norm minimization of the pseudo-inverse
Lemma 5 in Section 8.4 implies that the pseudo-inverse of a matrix with $k$ unit-norm columns has a minimal Schatten $p$-norm of $k^{1/p}$. We derive an inapproximability threshold of $k^{1/p}\sqrt{(1+\tfrac{1}{4k})}$ for $p>2$.
Theorem 11
It is NP hard to approximate the Schatten $p$-norm minimization of the pseudo-inverse within a factor of $\sqrt{1+\tfrac{1}{4k}}$ for $p>2$. Thus, unless P=NP, there is no PTAS for Schatten $p$-norm minimization of the pseudo-inverse.
Proof 11
The setup is the same as in the proof of Theorem 2. Assume that X3C instance is false, and assume that any two columns of $\boldsymbol{C}$ overlap. As in the proof of Theorem 10 conclude that $\|\boldsymbol{C}^{\dagger}\|_{F}\geq\sqrt{k+\tfrac{1}{4}}=k^{1/2}\sqrt{1+\tfrac{1}{4k}}$. For $p>2$, we interpret the Schatten $p$-norms as vector $p$-norms on the singular values, and apply the relation [26, (5.4.21)] between vector $p$-norms,
$\|\boldsymbol{C}^{\dagger}\|_{(p)}\geq k^{(\tfrac{1}{p}-\tfrac{1}{2})}\|\boldsymbol{C}^{\dagger}\|_{F}\geq k^{1/p}\sqrt{1+\tfrac{1}{4k}}.$
6 Condition numbers
We show that the minimization of two-norm, Frobenius norm, Schatten $p$-norm condition numbers, mixed and Schatten $p$-norm mixed condition numbers for $p>2$ is NP hard (Section 6.1), derive a new approximability threshold for 2-norm condition minimization (Section 6.2), and show that Frobenius norm condition number minimization (Section 6.3) and mixed condition number minimization (Section 6.4) do not admit a PTAS.
For $p\geq 1$, the $p$-norm condition number with regard to left inversion of a matrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$ is defined as
$\kappa_{(p)}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{(p)}\|\boldsymbol{C}^{\dagger}\|_{(p)}.$
Special cases include the two-norm condition number
$\kappa_{2}(\boldsymbol{C})=\|\boldsymbol{C}\|_{2}\|\boldsymbol{C}^{\dagger}\|_{2}=\sigma_{1}(\boldsymbol{C})/\sigma_{k}(\boldsymbol{C}),$
and the Frobenius norm condition number $\kappa_{F}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{F}\|\boldsymbol{C}^{\dagger}\|_{F}$.
The mixed condition number [15, Section 3] is $\kappa_{D}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{F}\|\boldsymbol{C}^{\dagger}\|_{2}$ and its extension to Schatten $p$-norms for $p>2$ is
$\kappa_{D,p}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{(p)}\|\boldsymbol{C}^{\dagger}\|_{2}.$
We want to find $k$ columns $\boldsymbol{C}$ of $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with minimal condition number, $\min_{\boldsymbol{C}}{\kappa_{\xi}(\boldsymbol{C})}$ for $\xi\in\{2,F,(p),D,(D,p)\}$.
6.1 Decision problems
Given a matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with unit-norm columns $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$; integer $1\leq k\leq\min\{\operatorname{\mathrm{rank}}(\boldsymbol{A}),m/3\}$; and parameter $b>0$.
  1. 1.
    Two-norm condition number minimization:
    Does $\boldsymbol{A}$ have a submatrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\kappa_{2}(\boldsymbol{C})\leq b$?
  2. 2.
    Frobenius-norm condition number minimization:
    Does $\boldsymbol{A}$ have a submatrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\kappa_{F}(\boldsymbol{C})\leq b$?
  3. 3.
    Schatten $p$-norm condition number minimization:
    Does $\boldsymbol{A}$ have a submatrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\kappa_{(p)}(\boldsymbol{C})\leq b$?
  4. 4.
    Mixed condition number minimization:
    Does $\boldsymbol{A}$ have a submatrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\kappa_{D}(\boldsymbol{C})\leq b$?
  5. 5.
    Schatten $p$-norm mixed condition number minimization
    Does $\boldsymbol{A}$ have a submatrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\kappa_{D,p}(\boldsymbol{C})\leq b$?
The NP hardness of two-norm condition number minimization was already established in [9, Theorem 4].
Theorem 12
The decision problems of condition number minimization are NP hard: in the two-norm, Frobenius norm, Schatten $p$-norms for $p>2$, mixed norm, and Schatten $p$-mixed norms for $p>2$.
Proof 12
The proof of NP hardness is the same as that of Theorem 3, but uses Lemma 6 instead.
6.2 No PTAS for two-norm condition number minimization
Lemma 6 in Section 8.5 implies that the minimal two-norm condition number is equal to 1. We derive an inapproximability threshold of $\sqrt{2}$.
Theorem 13
It is NP hard to approximate two-norm condition number minimization within a factor of $\sqrt{2}$. Thus, unless P=NP, there is no PTAS for two-norm condition number minimization.
Proof 13
The setup is the same as in the proof of Theorem 2. We derive a lower bound for the condition numbers for the reduced instance of X3C when it is false. Lemma 7 in Section 8.5 implies that the condition numbers decrease with the removal of columns. Hence the smallest condition numbers are found by identifying two columns for a false instance of X3C.
When the sets share a single element, the nonzero elements (6) of the relevant two columns $\hat{\boldsymbol{C}}$ have singular values $2/\sqrt{3}$ and $\sqrt{2/3}$. When the sets share two elements, the nonzero elements (7) of the relevant two columns $\hat{\boldsymbol{C}}$ have singular values $\sqrt{5/3}$ and $1/\sqrt{3}$. Between the two, (6) achieves the smallest condition number $(2/\sqrt{3})/(\sqrt{2/3})=\sqrt{2}$, which implies $\kappa_{2}(\boldsymbol{C})\geq\sqrt{2}$.
Theorem 13 derives a PTAS inapproximability threshold of $\sqrt{2}$ which differs from the one in [9, Theorem 8], which is $(2^{2k-3}/3^{k-2})^{\frac{1}{2(k-1)}}$.
6.3 No PTAS for Frobenius norm condition number minimization
Lemma 6 in Section 8.5 implies that a matrix with $k$ unit-norm columns has a minimal Frobenius norm condition number of $k$. We derive an inapproximability threshold of $\sqrt{k\left(k+\tfrac{1}{4}\right)}$.
Theorem 14
It is NP hard to approximate Frobenius norm condition number minimization within a factor of $\sqrt{1+\tfrac{1}{4k}}$. Thus, unless P=NP, there is no PTAS for Frobenius norm condition number minimization.
Proof 14
From Theorem 10 follows $\|\boldsymbol{C}^{\dagger}\|_{F}\geq\sqrt{k+\tfrac{1}{4}}$. Regardless of overlaps, the columns of $\boldsymbol{C}$ have unit norm, so (10) in Section 8.1 implies $\|\boldsymbol{C}\|_{F}=\sqrt{k}$. Thus
$\kappa_{F}(\boldsymbol{C})\geq\sqrt{k(k+\tfrac{1}{4})}=k\sqrt{1+\tfrac{1}{4k}}.$
6.4 No PTAS for mixed condition number minimization
Lemma 6 in Section 8.5 implies that a matrix with $k$ unit-norm columns has a mixed condition number with a minimal value of $\sqrt{k}$. We derive an approximability threshold of $\sqrt{\tfrac{3}{2}k}$.
Theorem 15
It is NP hard to approximate mixed condition number minimization within a factor of $\sqrt{3/2}$. Thus, unless P=NP, there is no PTAS for mixed condition number minimization.
Proof 15
The setup is the same as in the proof of Theorem 2. Assume that X3C instance is false, and that any two columns of $\boldsymbol{C}$ overlap by at least one element. Theorem 9 implies
$\kappa_{D}(\boldsymbol{C})^{2}=\|\boldsymbol{C}\|_{F}^{2}\|\boldsymbol{C}^{\dagger}\|_{2}^{2}\geq\frac{3}{2}\,\|\boldsymbol{C}\|_{F}^{2}.$
Regardless of overlaps, the columns of $\boldsymbol{C}$ have unit norm, so (10) in Section 8.1 implies $\|\boldsymbol{C}\|_{F}^{2}=k$. Thus $\kappa_{D}(\boldsymbol{C})^{2}\geq\frac{3}{2}k$.
7 Stable rank
We show that maximization of the stable rank and the $p$-stable rank for $p>2$ is NP hard (Section 7.1), and show that stable rank maximization has no PTAS (Section 7.2).
The stable rank of a nonzero matrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ is defined as
$\operatorname{\mathrm{sr}}(\boldsymbol{C})\equiv{\|\boldsymbol{C}\|_{F}^{2}}/{\|\boldsymbol{C}\|_{2}^{2}},$
and $\operatorname{\mathrm{sr}}(\boldsymbol{0}_{m\times k})=0$. The extension to Schatten $p$-norms for $p>2$ [27] is
$\operatorname{\mathrm{sr}}_{(p)}(\boldsymbol{C})\equiv{\|\boldsymbol{C}\|_{(p)}^{p}}/{\|\boldsymbol{C}\|_{2}^{p}},$
and $\operatorname{\mathrm{sr}}_{(p)}(\boldsymbol{0}_{m\times k})=0$. Obviously, $\operatorname{\mathrm{sr}}_{(2)}(\boldsymbol{A})=\operatorname{\mathrm{sr}}(\boldsymbol{A})$.
We want to find $k$ columns $\boldsymbol{C}$ of $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ that maximize the $p$-stable rank $\max_{\boldsymbol{C}}{\operatorname{\mathrm{sr}}_{(p)}(\boldsymbol{C})}$ for $p\geq 2$.
7.1 Decision problems
Given a matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with unit-norm columns $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$; $1\leq k\leq\min\{\operatorname{\mathrm{rank}}(\boldsymbol{A}),m/3\}$; and parameter $b>0$.
  1. 1.
    Stable rank maximization:
    Does $\boldsymbol{A}$ have a submatrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{sr}}(\boldsymbol{C})\geq b$?
  2. 2.
    $p$-stable rank maximization:
    Does $\boldsymbol{A}$ have a submatrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{sr}}_{p}(\boldsymbol{C})\geq b$?
Theorem 16
The decision problems of stable rank maximization and $p$-stable rank maximization for $p>2$ are NP hard.
Proof 16
The proof is the same as that of Theorem 3 but uses Lemma 8 in Section 8.6 instead.
7.2 No PTAS for stable rank maximization
Lemma 8 in Section 8.6 implies that, for a matrix with $k$ unit-norm columns, the maximal value of the stable rank equals $k$. We derive an inapproximability threshold of $\tfrac{3}{4}k$.
Theorem 17
It is NP hard to approximate stable rank maximization within a factor of $3/4$. Thus, unless P=NP, there is no PTAS for stable rank maximization.
Proof 17
The setup is the same as in the proof of Theorem 2. Assume that X3C instance is false, in which case Theorem 7 implies $\|\boldsymbol{C}\|_{2}\geq 2/\sqrt{3}$. Hence
$\operatorname{\mathrm{sr}}(\boldsymbol{C})=\frac{\|\boldsymbol{C}\|_{F}^{2}}{\|\boldsymbol{C}\|_{2}^{2}}\leq\frac{3}{4}\|\boldsymbol{C}\|_{F}^{2}=\frac{3}{4}k,$
because, regardless of overlaps, the columns of $\boldsymbol{C}$ have unit norm, so (10) in Section 8.1 implies $\|\boldsymbol{C}\|_{F}^{2}=k$.
8 Optimal values for subset selection criteria
To express the optimization versions of the subset selection criteria as decision problems, we derive optimal values for: maximal volume and S-optimality (Section 8.1), maximal relative volume (Section 8.2), minimal two-norm and Schatten $p$-norms (Section 8.3), minimal pseudo-inverse norms (Section 8.4), minimal condition numbers (Section 8.5), and maximal stable ranks (Section 8.6).
8.1 Maximal volume and S-optimality
We show that for matrices with unit-norm columns, the volume and S-optimality are bounded above by 1; and that only matrices with orthonormal columns have a maximal volume and S-optimality equal to 1.
Lemma 1
Let $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$ and $\|\boldsymbol{C}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq k$. Then each of the following inequalities holds with equality if and only if $\boldsymbol{C}$ has orthonormal columns.
  1. 1.
    Volume: $\operatorname{\mathrm{vol}}(\boldsymbol{C})\leq 1$.
  2. 2.
    S-optimality: $\operatorname{\mathrm{s_{opt}}}(\boldsymbol{C})\leq 1$.
Proof 18
The assumptions imply
$\|\boldsymbol{C}\|_{F}^{2}=\sum_{j=1}^{k}\|\boldsymbol{C}\boldsymbol{e}_{j}\|_{2}^{2}=k.$ (10)
We tackle each criterion, in turn.
  1. 1.
    Volume maximization: From (10) and the relation between the arithmetic and geometric means follows
    $\begin{split}1=\frac{\|\boldsymbol{C}\|_{F}^{2}}{k}=\frac{\sigma_{1}(\boldsymbol{C})^{2}+\cdots+\sigma_{k}(\boldsymbol{C})^{2}}{k}&\geq\sqrt[k]{\sigma_{1}(\boldsymbol{C})^{2}\cdots\sigma_{k}(\boldsymbol{C})^{2}}\\ &=\operatorname{\mathrm{vol}}(\boldsymbol{C})^{2/k}.\end{split}$ (11)
    Thus $\operatorname{\mathrm{vol}}(\boldsymbol{C})\leq 1$, which shows the inequality. Alternatively this also follows from Hadamard’s inequality [26, Corollary 7.8.3] and the fact that the columns have unit two-norm.
    As for the equality, if $\boldsymbol{C}$ has orthonormal columns, then all singular values $\sigma_{j}(\boldsymbol{C})=1$, $1\leq j\leq k$. Hence $\operatorname{\mathrm{vol}}(\boldsymbol{C})=1$.
    Conversely, if $\operatorname{\mathrm{vol}}(\boldsymbol{C})=1$, then $\operatorname{\mathrm{vol}}(\boldsymbol{C})^{2/k}=1$, and (11) implies that the arithmetic and geometric means are the same. This can only happen if all $\sigma_{j}(\boldsymbol{C})$ are equal. Thus, $\sigma_{j}(\boldsymbol{C})=1$, $1\leq j\leq k$, and $\boldsymbol{C}$ has orthonormal columns.
  2. 2.
    S-optimality maximization: The inequality follows from Hadamard’s inequality [26, Theorem 7.8.1],
    $\operatorname{\mathrm{vol}}(\boldsymbol{C})^{2}=\det(\boldsymbol{C}^{T}\boldsymbol{C})\leq\prod_{i=1}^{k}{(\boldsymbol{C}^{T}\boldsymbol{C})_{ii}}=\prod_{i=1}^{k}{\|\boldsymbol{C}\boldsymbol{e}_{i}\|_{2}^{2}}.$
    The equality follows from the assumption that $\boldsymbol{C}$ has unit-norm columns, so that $\operatorname{\mathrm{s_{opt}}}(\boldsymbol{S})=\operatorname{\mathrm{vol}}(\boldsymbol{C})^{1/k}$, and from the previous part. A more complicated proof of the equality is presented in [39, Theorem 3.2].
8.2 Maximal relative volume
We show that the relative volume is bounded above by 1, and that only matrices with orthonormal columns have a relative volume equal to 1.
Lemma 2
Let $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$. Then $0<\operatorname{\mathrm{rvol}}(\boldsymbol{C})\leq 1$.
Furthermore, $\operatorname{\mathrm{rvol}}(\boldsymbol{C})=1$ if and only if the columns of $\boldsymbol{C}$ are orthonormal.
Proof 19
The inequality follows from the singular value ordering (1).
As for the equality, if $\boldsymbol{C}$ has orthonormal columns, then all singular values of $\boldsymbol{C}$ are equal to 1, and $\operatorname{\mathrm{rvol}}(\boldsymbol{C})=1$. Conversely, if $\operatorname{\mathrm{rvol}}(\boldsymbol{C})=1$ then
$\prod_{j=1}^{k}{\frac{\sigma_{j}(\boldsymbol{C})}{\sigma_{1}(\boldsymbol{C})}}=1.$ (12)
Since all factors $\sigma_{j}(\boldsymbol{C})/\sigma_{1}(\boldsymbol{C})\leq 1$, (12) can only hold if all factors $\sigma_{j}(\boldsymbol{C})/\sigma_{1}(\boldsymbol{C})=1$. Hence, all singular values of $\boldsymbol{C}$ are equal to 1, and $\boldsymbol{C}$ has orthonormal columns.
We show that removal of columns can only increase the relative volume.
Lemma 3
Let $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$, and let $\boldsymbol{C}_{\ell}\in\mathbb{R}^{m\times\ell}$ be a column submatrix of $\boldsymbol{C}$, $1\leq\ell<k$. Then
$\operatorname{\mathrm{rvol}}(\boldsymbol{C})\leq\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{\ell})\leq\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{1})=1.$
Proof 20
Since $\boldsymbol{C}$ has full column rank, all columns are non-zero, and any non-zero column $\boldsymbol{C}_{1}$ attains the largest relative volume of 1.
For $\ell\geq 2$, singular value interlacing (2) implies
$\sigma_{k-\ell+j}(\boldsymbol{C})\leq\sigma_{j}(\boldsymbol{C}_{\ell})\leq\sigma_{j}(\boldsymbol{C}),\qquad 1\leq j\leq\ell.$
Since all factors $\sigma_{j}(\boldsymbol{C})/\sigma_{1}(\boldsymbol{C})\leq 1$,
$\operatorname{\mathrm{rvol}}(\boldsymbol{C})=\prod_{j=1}^{k}{\frac{\sigma_{j}(\boldsymbol{C})}{\sigma_{1}(\boldsymbol{C})}}\leq\prod_{j=1}^{\ell}{\frac{\sigma_{k-\ell+j}(\boldsymbol{C})}{\sigma_{1}(\boldsymbol{C})}}\leq\prod_{j=1}^{\ell}{\frac{\sigma_{j}(\boldsymbol{C}_{\ell})}{\sigma_{1}(\boldsymbol{C}_{\ell})}}=\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{\ell}).$
8.3 Minimal two-norm and Schatten p-norms
We show that only matrices with $k$ orthonormal columns have a two-norm equal to 1, and a Schatten $p$-norm equal to $k^{1/p}$ for $p>2$. In the special case $p=2$, the equality $\|\boldsymbol{C}\|_{(2)}=\|\boldsymbol{C}\|_{F}=\sqrt{k}$ is always true for unit-norm columns, regardless of whether they are orthonormal.
Lemma 4
Let $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$ and $\|\boldsymbol{C}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq k$. Then each of the following inequalities holds with equality if and only if $\boldsymbol{C}$ has orthonormal columns.
  • 1.
    Two-norm: $\|\boldsymbol{C}\|_{2}\geq 1$
  • 2.
    Schatten $p$-norm: $\|\boldsymbol{C}\|_{(p)}\geq k^{1/p}$ for $p>2$.
Proof 21
We tackle the two cases separately.
  1. 1.
    Two-norm: The inequality follows from the definition of the norm
    $\|\boldsymbol{C}\|_{2}\geq\max_{1\leq i\leq k}\|\boldsymbol{C}\boldsymbol{e}_{i}\|_{2}=1.$
    As for the equality, if $\boldsymbol{C}$ has orthonormal columns, then $\|\boldsymbol{C}\|_{2}=\sigma_{1}(\boldsymbol{C})=1$.
    Conversely, if $\|\boldsymbol{C}\|_{2}=1$, then (10) in Section 8.1 and the singular value ordering imply $\sigma_{1}(\boldsymbol{C})=\dots=\sigma_{k}(\boldsymbol{C})=1$. That is, $\boldsymbol{C}$ has orthonormal columns.
  2. 2.
    Schatten $p$-norm for $p>2$: The norm inequalities [26, Equation (5.4.21)] imply
    $\sqrt{k}=\|\boldsymbol{C}\|_{F}\leq k^{(\tfrac{1}{2}-\tfrac{1}{p})}\|\boldsymbol{C}\|_{(p)},\qquad p>2.$
    Rearranging gives the inequality $\|\boldsymbol{C}\|_{(p)}\geq k^{1/p}$ for $p>2$.
    As for the equality, we use the same idea as in the proof of [19, Proposition 3.1]. If $\boldsymbol{C}$ has orthonormal columns, then all the singular values of $\boldsymbol{C}$ are equal to $1$ and $\|\boldsymbol{C}\|_{(p)}=k^{1/p}$.
    Conversely, if $\|\boldsymbol{C}\|_{(p)}=k^{1/p}$, then (10) in Section 8.1 implies
    $\sum_{j=1}^{k}\left(\frac{2}{p}\sigma_{j}(\boldsymbol{C})^{p}-\sigma_{j}(\boldsymbol{C})^{2}+(1-\tfrac{2}{p})\right)=0.$ (13)
    This sum has the form
    $\sum_{j=1}^{k}g_{p}(\sigma_{j}(\boldsymbol{C}))=0\qquad\text{where}\quad g_{p}(x)\equiv\tfrac{2x^{p}}{p}-x^{2}+(1-\tfrac{2}{p}).$
    For $x>0$, $g_{p}(x)$ has the unique global minimizer $x_{*}=1$ with $g_{p}(x_{*})=0$. Since the singular values are positive, $\sum_{j=1}^{k}g(\sigma_{j}(\boldsymbol{C}))$ is a sum of nonnegative summands. Thus equality in (13) is achieved only if $\sigma_{1}(\boldsymbol{C})=\dots=\sigma_{k}(\boldsymbol{C})=1$, implying that $\boldsymbol{C}$ has orthonormal columns.
Below is a justification for limiting Schatten $p$-norms to the case $p\geq 2$: The lower bound $\|\boldsymbol{C}\|_{(p)}\geq k^{1/p}$ for $p>2$ turns into an upper bound for $1\leq p<2$.
Remark 8.1
Let $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$ and $\|\boldsymbol{C}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq k$. For $1\leq p<2$, [26, (5.4.21)] implies
$\sqrt{k}\leq\|\boldsymbol{C}\|_{(p)}\leq k^{(\tfrac{1}{p}-\tfrac{1}{2})}\sqrt{k}=k^{1/p}$
and $\|\boldsymbol{C}\|_{(p)}=k^{1/p}$ if and only if $\boldsymbol{C}$ has orthonormal columns.
8.4 Minimal pseudo-inverse norms
We show that matrices with $k$ unit-norm columns have a pseudo-inverse whose two-norm is bounded below by $1$, and whose $p$-norm is bounded below by $k^{1/p}$ for $p\geq 2$; and that only matrices with $k$ orthonormal columns have pseudo-inverses with a two-norm equal to 1, and a Schatten $p$-norm equal to $k^{1/p}$ for $p>2$.
Lemma 5
Let $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$ and $\|\boldsymbol{C}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq k$. Then each of the following inequalities holds with equality if and only if $\boldsymbol{C}$ has orthonormal columns.
  • 1.
    Two-norm of pseudo-inverse: $\|\boldsymbol{C}^{\dagger}\|_{2}\geq 1$
  • 2.
    Frobenius-norm of pseudo-inverse: $\|\boldsymbol{C}^{\dagger}\|_{F}\geq\sqrt{k}$
  • 3.
    Schatten $p$-norm of pseudo-inverse: $\|\boldsymbol{C}^{\dagger}\|_{(p)}\geq k^{1/p}$.
Proof 22
We tackle each case separately.
  • 1.
    Two-norm of the pseudo-inverse: From (10) in Section 8.1 follows
    $k=\|\boldsymbol{C}\|_{F}^{2}=\sum_{j=1}^{k}{\sigma_{j}(\boldsymbol{C})^{2}}\geq k\sigma_{k}(\boldsymbol{C})^{2}$
    Thus $\sigma_{k}(\boldsymbol{C})\leq 1$ and $\|\boldsymbol{C}^{\dagger}\|_{2}=1/\sigma_{k}(\boldsymbol{C})\geq 1$, which shows the inequality.
    As for the equality, if $\boldsymbol{C}$ has orthonormal columns, then $\boldsymbol{C}^{\dagger}=\boldsymbol{C}^{T}$, hence $\|\boldsymbol{C}^{\dagger}\|_{2}=1$.
    Conversely, if $\|\boldsymbol{C}^{\dagger}\|_{2}=1$, then the singular value ordering $\sigma_{1}(\boldsymbol{C})\geq\cdots\geq\sigma_{k}(\boldsymbol{C})=1$ and (10) in Section 8.1 imply $\sigma_{j}(\boldsymbol{C})=1$, $1\leq j\leq k$. Thus, $\boldsymbol{C}$ has orthonormal columns.
  • 2.
    Frobenius norm of the pseudo-inverse: This is a special case of the Schatten $p$-norm for $p=2$.
  • 3.
    Schatten p-norm of the pseudo-inverse: From
    $k=\|\boldsymbol{C}\|_{F}^{2}=\sum_{j=1}^{k}\sigma_{j}(\boldsymbol{C})^{2}\leq k\sigma_{1}(\boldsymbol{C})^{2}$
    follows $\sigma_{1}(\boldsymbol{C})\geq 1$. Therefore,
    $\|\boldsymbol{C}^{\dagger}\|_{(p)}\geq\left(\frac{k}{\sigma_{1}(\boldsymbol{C})^{p}}\right)^{1/p}\geq k^{1/p}.$
    As for the equality, if $\boldsymbol{C}$ has orthornormal columns, then $\sigma_{j}(\boldsymbol{C})=1$, $1\leq j\leq k$. Hence, $\|\boldsymbol{C}^{\dagger}\|_{(p)}=k^{1/p}$.
    Conversely, if $\|\boldsymbol{C}^{\dagger}\|_{(p)}=k^{1/p}$, then (10) in Section 8.1 implies
    $0=\|\boldsymbol{C}\|_{F}^{2}+\frac{2}{p}\|\boldsymbol{C}^{\dagger}\|_{(p)}^{p}-k(1+\tfrac{2}{p})=\sum_{j=1}^{k}(\sigma_{j}(\boldsymbol{C})^{2}+\tfrac{2}{p\sigma_{j}(\boldsymbol{C})^{p}}-(1+\tfrac{2}{p})).$ (14)
    Each summand equals $g(\sigma_{j}(\boldsymbol{C}))$ for $g(x)\equiv x^{2}+\tfrac{2}{px^{p}}-(1+\tfrac{2}{p})$. For $x>0$, the function $g(x)$ is nonnegative with global minimizer $x_{*}=1$ and $g(x_{*})=0$. Thus, equality in (14) is achieved if and only if $\sigma_{j}(\boldsymbol{C})=1$, $1\leq j\leq k$, and $\boldsymbol{C}$ has orthonormal columns.
8.5 Minimal condition numbers
We derive lower bounds for condition numbers of matrices with unit-norm columns, and show that the lower bounds are attained only by matrices with orthonormal columns.
Lemma 6
Let $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$ and $\|\boldsymbol{C}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq k$. Then each of the following inequalities holds with equality if and only if $\boldsymbol{C}$ has orthonormal columns.
  • 1.
    Two-norm condition number: $\kappa_{2}(\boldsymbol{C})\geq 1$
  • 2.
    Frobenius norm condition number: $\kappa_{F}(\boldsymbol{C})\geq k$
  • 3.
    Schatten $p$-norm condition number: $\kappa_{(p)}(\boldsymbol{C})\geq k^{2/p}$
  • 4.
    Mixed condition number: $\kappa_{D}(\boldsymbol{C})\geq\sqrt{k}$
  • 5.
    Schatten $p$-norm mixed condition number $\kappa_{D,p}(\boldsymbol{C})\geq k^{1/p}$
Proof 23
We tackle each case separately.
  1. 1.
    Two-norm condition number: The inequality follows from the singular value ordering (1).
    As for the equality, if $\boldsymbol{C}$ has orthonormal columns, then all its singular values are $1$ and $\kappa_{2}(\boldsymbol{C})=1$. Conversely, if $\kappa_{2}(\boldsymbol{C})=\sigma_{1}(\boldsymbol{C})/\sigma_{k}(\boldsymbol{C})=1$, then the singular value ordering (1) ensures that $\sigma_{j}(\boldsymbol{C})=1$, $1\leq j\leq k$, and $\boldsymbol{C}$ has orthonormal columns.
  2. 2.
    Frobenius norm condition number: The inequality follows from applying the Cauchy-Schwartz inequality to the vectors
    $\mathbf{x}\equiv\begin{bmatrix}\sigma_{1}(\boldsymbol{C})&\cdots&\sigma_{k}(\boldsymbol{C})\end{bmatrix}^{T},\qquad\boldsymbol{y}\equiv\begin{bmatrix}1/\sigma_{1}(\boldsymbol{C})&\cdots&1/\sigma_{k}(\boldsymbol{C})\end{bmatrix}^{T},$
    which shows
    $k=\mathbf{x}^{T}\boldsymbol{y}\leq\|\mathbf{x}\|_{2}\|\boldsymbol{y}\|_{2}=\left(\sum_{j=1}^{k}\sigma_{j}(\boldsymbol{C})^{2}\right)^{1/2}\left(\sum_{j=1}^{k}\sigma_{j}(\boldsymbol{C})^{-2}\right)^{1/2}=\kappa_{F}(\boldsymbol{C}).$
    As for the equality, if $\boldsymbol{C}$ has orthonormal columns, then $\kappa_{F}(\boldsymbol{C})=k$.
    Conversely, if $\kappa_{F}(\boldsymbol{C})=k$, then (10) in Section 8.1 implies $\|\boldsymbol{C}\|_{F}=\sqrt{k}$, hence $\|\boldsymbol{C}^{\dagger}\|_{F}^{2}=\sum_{j=1}^{k}\sigma_{j}(\boldsymbol{C})^{-2}=k$. As in the proof of Lemma 5, one shows $\sigma_{j}(\boldsymbol{C})=1$, $1\leq j\leq k$. Thus, $\boldsymbol{C}$ has orthonormal columns.
  3. 3.
    Schatten $p$-norm condition number for $p>2$: The inequality follows from applying the Cauchy-Schwartz inequality to the vectors
    $\mathbf{x}\equiv\begin{bmatrix}\sigma_{1}(\boldsymbol{C})^{p}&\cdots&\sigma_{k}(\boldsymbol{C})^{p}\end{bmatrix}^{T},\qquad\boldsymbol{y}\equiv\begin{bmatrix}1/\sigma_{1}(\boldsymbol{C})^{p}&\cdots&1/\sigma_{k}(\boldsymbol{C})^{p}\end{bmatrix}^{T},$
    which shows
    $k=\mathbf{x}^{T}\boldsymbol{y}\leq\|\mathbf{x}\|_{2}\|\boldsymbol{y}\|_{2}=\left(\sum_{j=1}^{k}\sigma_{j}(\boldsymbol{C})^{p}\right)^{1/2}\left(\sum_{j=1}^{k}\sigma_{j}(\boldsymbol{C})^{-p}\right)^{1/2}=\kappa_{p}(\boldsymbol{C})^{p/2}.$
    Equality occurs in the Cauchy-Schwartz inequality only if $\mathbf{x}$ is a multiple of $\boldsymbol{y}$. This ensures that all singular values are equal to $1$, thus $\boldsymbol{C}$ has orthonormal columns.
    Conversely, if $\boldsymbol{C}$ has orthonormal columns, then all the singular values are equal to $1$ and $\kappa_{(p)}(\boldsymbol{C})=k^{2/p}$.
  4. 4.
    Mixed condition number: The inequality follows from the definition of $\kappa_{D}(\boldsymbol{C})$. As for the equality, if $\boldsymbol{C}$ has orthonormal columns, then $\kappa_{D}(\boldsymbol{C})=\sqrt{k}$. Conversely, if $\kappa_{D}(\boldsymbol{C})=\sqrt{k}$, then (10) in Section 8.1 implies $\sigma_{k}(\boldsymbol{C})=1$. The singular value ordering $\sigma_{1}(\boldsymbol{C})\geq\cdots\geq\sigma_{k}(\boldsymbol{C})=1$ and (10) imply $\sigma_{j}(\boldsymbol{C})=1$, $1\leq j\leq k$. Thus $\boldsymbol{C}$ has orthonormal columns.
  5. 5.
    Schatten $p$-norm mixed condition number for $p>2$: The inequality follows from the definition of $\kappa_{D,p}(\boldsymbol{C})$. As for the equality, if $\boldsymbol{C}$ has orthonormal columns, then $\kappa_{D,p}(\boldsymbol{C})={k}^{1/p}$.
    Conversely, if $\kappa_{D,p}(\boldsymbol{C})={k}^{1/p}$, then $\sum_{j=1}^{{\color[rgb]{0,0,0}k}}\sigma_{j}(\boldsymbol{C})^{p}=k\sigma_{k}(\boldsymbol{C})^{p}$, so that
    $\sum_{j=1}^{{\color[rgb]{0,0,0}k}}(\sigma_{j}(\boldsymbol{C})^{p}-\sigma_{k}(\boldsymbol{C})^{p})=0.$
    The singular value ordering implies that all singular values are equal to $1$, thus $\boldsymbol{C}$ has orthonormal columns.
We show that the removal of columns can only increase the condition numbers.
Lemma 7
Let $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$, and let $\boldsymbol{C}_{\ell}\in\mathbb{R}^{m\times\ell}$ be a column submatrix of $\boldsymbol{C}$, $1\leq\ell<k$. Then
$\kappa_{\xi}(\boldsymbol{C})$ $\geq\kappa_{\xi}(\boldsymbol{C}_{\ell})\geq\kappa_{\xi}(\boldsymbol{C}_{1})=1,\qquad\xi\in\{2,F,(p),D,(D,p)\}.$
Proof 24
Since $\boldsymbol{C}$ has full column rank, all columns are non-zero, and any non-zero column $\boldsymbol{C}_{1}$ attains the minimal value of 1 for all the condition numbers.
  • 1.
    Two-norm condition number: For $\ell\geq 2$, singular value interlacing (2) implies $\sigma_{1}(\boldsymbol{C})\geq\sigma_{1}(\boldsymbol{C}_{\ell})$ and $\sigma_{k}(\boldsymbol{C})\leq\sigma_{\ell}(\boldsymbol{C}_{\ell})$, hence
    $\kappa_{2}(\boldsymbol{C})=\frac{\sigma_{1}(\boldsymbol{C})}{\sigma_{k}(\boldsymbol{C})}\geq\frac{\sigma_{1}(\boldsymbol{C}_{\ell})}{\sigma_{\ell}(\boldsymbol{C}_{\ell})}=\kappa_{2}(\boldsymbol{C}_{\ell}).$
  • 2.
    Frobenius norm and Schatten-$p$ norm condition numbers for $p>2$: We give the proof for $p\geq 2$, which includes the case of the Frobenius norm. For $\ell\geq 2$, singular value interlacing (2) implies
    $\sigma_{k-\ell+j}(\boldsymbol{C})\leq\sigma_{j}(\boldsymbol{C}_{\ell})\leq\sigma_{j}(\boldsymbol{C}),\qquad 1\leq j\leq\ell.$
    Thus
    $\|\boldsymbol{C}\|_{(p)}^{p}=\sum_{j=1}^{k}{\sigma_{j}(\boldsymbol{C})^{p}}\geq\sum_{j=1}^{\ell}{\sigma_{j}(\boldsymbol{C})^{p}}\geq\sum_{j=1}^{\ell}{\sigma_{j}(\boldsymbol{C}_{\ell})^{p}}=\|\boldsymbol{C}_{\ell}\|_{(p)}^{p}$
    and
    $\|\boldsymbol{C}^{\dagger}\|_{(p)}^{p}=\sum_{j=1}^{k}{\frac{1}{\sigma_{j}(\boldsymbol{C})^{p}}}\geq\sum_{j=k-\ell+1}^{\ell}{\frac{1}{\sigma_{j}(\boldsymbol{C})^{p}}}\geq\sum_{j=1}^{\ell}{\frac{1}{\sigma_{j}(\boldsymbol{C}_{\ell})^{p}}}=\|\boldsymbol{C}_{\ell}^{\dagger}\|_{(p)}^{p}.$
    Thus
    $\left(\kappa_{(p)}(\boldsymbol{C})\right)^{p}=\|\boldsymbol{C}\|_{(p)}^{p}\|\boldsymbol{C}^{\dagger}\|_{(p)}^{p}\geq\|\boldsymbol{C}_{\ell}\|_{(p)}^{p}\|\boldsymbol{C}_{\ell}^{\dagger}\|_{(p)}^{p}=\left((\kappa_{(p)}(\boldsymbol{C}_{\ell})\right)^{p}.$
  • 3.
    Mixed condition number and Schatten $p$-norm mixed condition numbers: The proofs are analogous to the ones above.
8.6 Maximal stable rank
We derive upper bounds for the stable rank and $p$-stable rank for $p>2$ of matrices with unit-norm columns, and show that the upper bounds are attained only by matrices with orthonormal columns.
Lemma 8
Let $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$ and $\|\boldsymbol{C}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq k$. Then each of the following inequalities holds with equality if and only if $\boldsymbol{C}$ has orthonormal columns.
  1. 1.
    Stable rank: $\operatorname{\mathrm{sr}}(\boldsymbol{C})\leq k$
  2. 2.
    $p$-stable rank: $\operatorname{\mathrm{sr}}_{(p)}(\boldsymbol{C})\leq k$
Proof 25
We tackle each part in turn.
  1. 1.
    Stable rank: The inequality follows from the definition.
    As for the equality, if $\boldsymbol{C}$ has orthonormal columns, then all $k$ singular values equal to $1$, and $\operatorname{\mathrm{sr}}(\boldsymbol{C})=k$. Conversely, if $k=\operatorname{\mathrm{sr}}(\boldsymbol{C})=\|\boldsymbol{C}\|_{F}^{2}/\|\boldsymbol{C}\|_{2}^{p}$, then the definition of stable rank implies
    $\sum_{j=1}^{k}{\sigma_{j}(\boldsymbol{C})^{2}}=\|\boldsymbol{C}\|_{F}^{2}=k\|\boldsymbol{C}\|_{2}^{2}=k\,\sigma_{1}(\boldsymbol{C})^{2}.$
    Hence $\sum_{j=1}^{k}{(\sigma_{1}(\boldsymbol{C})^{2}-\sigma_{j}(\boldsymbol{C})^{2})}=0$. From the singular value ordering (1) follows that this is a sum of non-negative summands, and every summand must be equal to zero. Hence, all singular values are the same. From (10) in Section 8.1 follows that they must be equal to one, so $\boldsymbol{C}$ has orthonormal columns.
  2. 2.
    $p$-stable rank for $p>2$: The definition of $p$-stable rank implies $\operatorname{\mathrm{sr}}_{(p)}(\boldsymbol{C})\leq k$.
    As for the equality, if $\boldsymbol{C}$ has orthonormal columns, then all $k$ singular values of $\boldsymbol{C}$ are equal to $1$, and $\operatorname{\mathrm{sr}}_{(p)}(\boldsymbol{C})=k$, see also [27, Example 2.1].
    Conversely, if $k=\operatorname{\mathrm{sr}}_{(p)}(\boldsymbol{C})=\|\boldsymbol{C}\|_{(p)}^{p}/\|\boldsymbol{C}\|_{2}^{p}$, then the definition of $p$-stable rank implies
    $\sum_{j=1}^{k}{\sigma_{j}(\boldsymbol{C})^{p}}=\|\boldsymbol{C}\|_{(p)}^{p}=k\|\boldsymbol{C}\|_{2}^{p}=k\,\sigma_{1}(\boldsymbol{C})^{p}.$
    Hence $\sum_{j=1}^{k}{(\sigma_{1}(\boldsymbol{C})^{p}-\sigma_{j}(\boldsymbol{C})^{p})}=0$. From the singular value ordering (1) follows that this is a sum of non-negative summands, and every summand must be equal to zero. Hence, all singular values of $\boldsymbol{C}$ are the same. From (10) in Section 8.1 follows that they must be equal to one, so $\boldsymbol{C}$ has orthonormal columns.
9 Expressions for partitioned pseudo-inverses
For partitioned full column-rank matrices, we derive expressions for the partitioned pseudo-inverses (Lemma 9), and lower bounds for the Frobenius norm of the partitioned pseudo-inverse (Lemma 10) and for Schatten $p$-norms for $p>2$ (Lemma 11).
Different versions of the expressions below have appeared in [3, Theorem 1], [11, Theorem 2] and [7, Lemma 3.3 and 3.4]. Here we present a short, self-contained proof.
Lemma 9
Let $\boldsymbol{C}=\begin{bmatrix}\boldsymbol{C}_{1}&\boldsymbol{C}_{2}\end{bmatrix}\in{\mathbb{R}}^{m\times n}$ with $\boldsymbol{C}_{1}\in\mathbb{R}^{m\times k}$ and $\boldsymbol{C}_{2}\in\mathbb{R}^{m\times(n-k)}$ for some $1\leq k<n$. If $\operatorname{\mathrm{rank}}(\boldsymbol{C})=n$, then
$\boldsymbol{C}^{\dagger}=\begin{bmatrix}\boldsymbol{M}_{1}^{\dagger}\\ \boldsymbol{M}_{2}^{\dagger}\end{bmatrix}\in\mathbb{R}^{n\times m},\qquad\text{where}\quad\boldsymbol{M}_{1}\equiv\boldsymbol{P}_{2}\boldsymbol{C}_{1},\quad\boldsymbol{M}_{2}\equiv\boldsymbol{P}_{1}\boldsymbol{C}_{2}$
and
$\boldsymbol{P}_{1}\equiv\boldsymbol{I}-\boldsymbol{C}_{1}\boldsymbol{C}_{1}^{\dagger},\qquad\boldsymbol{P}_{2}\equiv\boldsymbol{I}-\boldsymbol{C}_{2}\boldsymbol{C}_{2}^{\dagger}$
are $m\times m$ orthogonal projectors onto $\operatorname{\mathrm{range}}(\boldsymbol{C}_{1})^{\perp}$ and $\operatorname{\mathrm{range}}(\boldsymbol{C}_{2})^{\perp}$, respectively.
Proof 26
Since $\boldsymbol{C}$ has full column rank, $\boldsymbol{C}^{\dagger}=(\boldsymbol{C}^{T}\boldsymbol{C})^{-1}\boldsymbol{C}^{T}$, and
$\boldsymbol{C}^{T}\boldsymbol{C}=\begin{bmatrix}\boldsymbol{C}_{1}^{T}\boldsymbol{C}_{1}&\boldsymbol{C}_{1}^{T}\boldsymbol{C}_{2}\\ \boldsymbol{C}_{2}^{T}\boldsymbol{C}_{1}&\boldsymbol{C}_{2}^{T}\boldsymbol{C}_{2}\end{bmatrix}$.
Furthermore, the column submatrices $\boldsymbol{C}_{1}$ and $\boldsymbol{C}_{2}$ also have full column rank, so that $\boldsymbol{C}_{1}^{T}\boldsymbol{C}_{1}\in\mathbb{R}^{k\times k}$ and $\boldsymbol{C}_{2}^{T}\boldsymbol{C}_{2}\in\mathbb{R}^{(n-k)\times(n-k)}$ are nonsingular.
From [12, (4)] follows
$(\boldsymbol{C}^{T}\boldsymbol{C})^{-1}=\begin{bmatrix}\boldsymbol{S}_{1}^{-1}&\\ &\boldsymbol{S}_{2}^{-1}\end{bmatrix}\begin{bmatrix}\boldsymbol{I}&-\boldsymbol{C}_{1}^{T}\boldsymbol{C}_{2}(\boldsymbol{C}_{2}^{T}\boldsymbol{C}_{2})^{-1}\\ -\boldsymbol{C}_{2}^{T}\boldsymbol{C}_{1}(\boldsymbol{C}_{1}^{T}\boldsymbol{C}_{1})^{-1}&\boldsymbol{I}\end{bmatrix}$,
where
$\begin{split}\boldsymbol{S}_{1}&\equiv\boldsymbol{C}_{1}^{T}\boldsymbol{C}_{1}-\boldsymbol{C}_{1}^{T}\boldsymbol{C}_{2}\boldsymbol{C}_{2}^{\dagger}\boldsymbol{C}_{1}=\boldsymbol{C}_{1}^{T}\>(\boldsymbol{I}-\boldsymbol{C}_{2}\boldsymbol{C}_{2}^{\dagger})\>\boldsymbol{C}_{1}=\boldsymbol{C}_{1}^{T}\boldsymbol{P}_{2}\boldsymbol{C}_{1}=\boldsymbol{M}_{1}^{T}\boldsymbol{M}_{1}\\ \boldsymbol{S}_{2}&\equiv\boldsymbol{C}_{2}^{T}\boldsymbol{C}_{2}-\boldsymbol{C}_{2}^{T}\boldsymbol{C}_{1}\boldsymbol{C}_{1}^{\dagger}\boldsymbol{C}_{2}=\boldsymbol{M}_{2}^{T}\boldsymbol{M}_{2}.\end{split}$ (15)
The nonsingularity of the Schur complements $\boldsymbol{S}_{1}$ and $\boldsymbol{S}_{2}$ follows from the nonsingularity of $\boldsymbol{C}^{T}\boldsymbol{C}$, $\boldsymbol{C}_{1}^{T}\boldsymbol{C}_{1}$ and $\boldsymbol{C}_{2}^{T}\boldsymbol{C}_{2}$, see (8) and [12, Section 1].
Multiplying $(\boldsymbol{C}^{T}\boldsymbol{C})^{-1}$ by $\boldsymbol{C}^{T}$ on the right gives
$\boldsymbol{C}^{\dagger}=\begin{bmatrix}\boldsymbol{S}_{1}^{-1}(\boldsymbol{C}_{1}^{T}-\boldsymbol{C}_{1}^{T}\boldsymbol{C}_{2}\boldsymbol{C}_{2}^{\dagger})\\ \boldsymbol{S}_{2}^{-1}(-\boldsymbol{C}_{2}^{T}\boldsymbol{C}_{1}\boldsymbol{C}_{1}^{\dagger}+\boldsymbol{C}_{2}^{T})\end{bmatrix}=\begin{bmatrix}(\boldsymbol{M}_{1}^{T}\boldsymbol{M}_{1})^{-1}\boldsymbol{C}_{1}^{T}\boldsymbol{P}_{2}\\ (\boldsymbol{M}_{2}^{T}\boldsymbol{M}_{2})^{-1}\boldsymbol{C}_{2}^{T}\boldsymbol{P}_{1}\end{bmatrix}=\begin{bmatrix}\boldsymbol{M}_{1}^{\dagger}\\ \boldsymbol{M}_{2}^{\dagger}\end{bmatrix}$. (16)
Below is a lower bound on the Frobenius norm in terms of the partitioned inverses.
Lemma 10
Let $\boldsymbol{C}=\begin{bmatrix}\boldsymbol{C}_{1}&\boldsymbol{C}_{2}\end{bmatrix}\in{\mathbb{R}}^{m\times n}$ with $\boldsymbol{C}_{1}\in\mathbb{R}^{m\times k}$ and $\boldsymbol{C}_{2}\in\mathbb{R}^{m\times(n-k)}$ for some $1\leq k<n$. If $\operatorname{\mathrm{rank}}(\boldsymbol{C})=n$, then
$\|\boldsymbol{C}^{\dagger}\|_{F}^{2}\geq\|\boldsymbol{C}_{1}^{\dagger}\|_{F}^{2}+\|\boldsymbol{C}_{2}^{\dagger}\|_{F}^{2}.$
Proof 27
Lemma 9 and (16) imply
$\boldsymbol{C}^{\dagger}=\begin{bmatrix}\boldsymbol{M}_{1}^{\dagger}\\ \boldsymbol{M}_{2}^{\dagger}\end{bmatrix},\qquad\text{where}\quad\boldsymbol{M}_{1}\equiv\boldsymbol{P}_{2}\boldsymbol{C}_{2},\quad\boldsymbol{M}_{2}\equiv\boldsymbol{P}_{1}\boldsymbol{C}_{2}$
and $\boldsymbol{P}_{1}$ and $\boldsymbol{P}_{2}$ are orthogonal projectors. Since $\|\boldsymbol{P}_{1}\|_{2}=\|\boldsymbol{P}_{2}\|_{2}=1$, the product inequalities for singular values [26, Problem 7.3.P16] imply
$\sigma_{i}(\boldsymbol{M}_{1})=\sigma_{i}(\boldsymbol{P}_{2}\boldsymbol{C}_{1})\leq\|\boldsymbol{P}_{2}\|_{2}\,\sigma_{i}(\boldsymbol{C}_{1})=\sigma_{i}(\boldsymbol{C}_{1}),\qquad 1\leq i\leq k.$
Because both $\boldsymbol{M}_{1}$ and $\boldsymbol{C}_{1}$ have full rank, we can invert the singular values
$1/\sigma_{i}(\boldsymbol{C}_{1})\leq 1/\sigma_{i}(\boldsymbol{M}_{1}),\qquad 1\leq i\leq k.$ (17)
Therefore $\|\boldsymbol{M}_{1}^{\dagger}\|_{F}\geq\|\boldsymbol{C}_{1}^{\dagger}\|_{F}$. This is a special case of [32, Theorem 3.2]. The proof for $\|\boldsymbol{M}_{2}^{\dagger}\|_{F}\geq\|\boldsymbol{C}_{2}^{\dagger}\|_{F}$ is analogous. Thus
$\|\boldsymbol{C}^{\dagger}\|_{F}^{2}=\|\boldsymbol{M}_{1}^{\dagger}\|_{F}^{2}+\|\boldsymbol{M}_{2}^{\dagger}\|_{F}^{2}\geq\|\boldsymbol{C}_{1}^{\dagger}\|_{F}^{2}+\|\boldsymbol{C}_{2}^{\dagger}\|_{F}^{2}.$
Below is a lower bound on the Schatten $p$-norms in terms of the partitioned inverses.
Lemma 11
Let $\boldsymbol{C}=\begin{bmatrix}\boldsymbol{C}_{1}&\boldsymbol{C}_{2}\end{bmatrix}\in{\mathbb{R}}^{m\times n}$ have $\operatorname{\mathrm{rank}}(\boldsymbol{C})=n$. Then for $p>2$
$\|\boldsymbol{C}^{\dagger}\|_{(p)}^{2}\geq\|\boldsymbol{C}_{1}^{\dagger}\|_{(p)}^{2}+\|\boldsymbol{C}_{2}^{\dagger}\|_{(p)}^{2}.$
Proof 28
From (17), follows $\sigma_{j}(\boldsymbol{C}_{1}^{\dagger})^{2}\leq\sigma_{j}(\boldsymbol{M}_{1}^{\dagger})^{2}$ for $1\leq j\leq k$, or $\|\boldsymbol{C}_{1}^{\dagger}\|_{(p)}^{2}\leq\|\boldsymbol{M}_{1}^{\dagger}\|_{(p)}^{2}$, with an analogous relation between $\boldsymbol{C}_{2}$ and $\boldsymbol{M}_{2}$. Lemma 9 implies
$\boldsymbol{C}^{\dagger}(\boldsymbol{C}^{\dagger})^{T}=\begin{bmatrix}\boldsymbol{M}_{1}^{\dagger}(\boldsymbol{M}_{1}^{\dagger})^{T}&\boldsymbol{M}_{1}^{\dagger}(\boldsymbol{M}_{2}^{\dagger})^{T}\\ \boldsymbol{M}_{2}^{\dagger}(\boldsymbol{M}_{1}^{\dagger})^{T}&\boldsymbol{M}_{2}^{\dagger}(\boldsymbol{M}_{2}^{\dagger})^{T}\end{bmatrix}$.
Applying the pinching inequality [4, Equation (IV.52)] gives
$\|\boldsymbol{C}^{\dagger}\|_{(p)}^{2}=\|\boldsymbol{C}^{\dagger}(\boldsymbol{C}^{\dagger})^{T}\|_{(p/2)}$ ${\geq\|\operatorname{\mathrm{diag}}\begin{pmatrix}\boldsymbol{M}_{1}^{\dagger}(\boldsymbol{M}_{1}^{\dagger})^{T}&\boldsymbol{M}_{2}^{\dagger}(\boldsymbol{M}_{2}^{\dagger})^{T}\end{pmatrix}\|_{(p/2)}}$
$=\>\|\boldsymbol{M}_{1}^{\dagger}(\boldsymbol{M}_{1}^{\dagger})^{T}\|_{(p/2)}+\|\boldsymbol{M}_{2}^{\dagger}(\boldsymbol{M}_{2}^{\dagger})^{T}\|_{(p/2)}$
$\geq\>\|\boldsymbol{C}_{1}^{\dagger}(\boldsymbol{C}_{1}^{\dagger})^{T}\|_{(p/2)}+\|\boldsymbol{C}_{2}^{\dagger}(\boldsymbol{C}_{2}^{\dagger})^{T}\|_{(p/2)}$
$=\>\|\boldsymbol{C}_{1}^{\dagger}\|_{(p)}^{2}+\|\boldsymbol{C}_{2}^{\dagger}\|_{(p)}^{2}.$
Acknowledgments
We are most grateful to three anonymous reviewers for their constructive feedback in regard to exposition and literature references.